### v1

# R-PINN: Recovery-type a-posteriori estimator enhanced adaptive PINN

Abstract
In recent years, with the advancements in machine learning and neural networks, algorithms using physics-informed neural networks (PINNs) to solve PDEs have gained widespread applications. While these algorithms are well-suited for a wide range of equations, they often exhibit suboptimal performance when applied to equations with large local gradients, resulting in substantial localized errors. To address this issue, this paper proposes an adaptive PINN algorithm designed to improve accuracy in such cases. The core idea of the algorithm is to adaptively adjust the distribution of collocation points based on the recovery-type a-posterior error of the current numerical solution, enabling a better approximation of the true solution. This approach is inspired by the adaptive finite element method. By combining the recovery-type a-posteriori estimator, a gradient-recovery estimator commonly used in the adaptive finite element method (FEM) with PINNs, we introduce the Recovery-type a-posteriori estimator enhanced adaptive PINN (R-PINN) and compare its performance with a typical adaptive PINN algorithm, FI-PINN. Our results demonstrate that R-PINN achieves faster convergence with fewer adaptive points and significantly outperforms in the cases with multiple regions of large errors than FI-PINN. Notably, our method is a hybrid numerical approach for solving partial differential equations, integrating adaptive FEM with PINNs.
keywords:
Adaptive PINN , Recovery-type a-posteriori estimator , Machine learning , Adaptive FEM
1 Introduction
The Physics-informed neural networks (PINNs) algorithm is a data-driven approach for solving partial differential equations (PDEs) based on deep learning frameworks [1]. Due to its distinct methodology, which diverges significantly from traditional PDE numerical techniques, and by leveraging the superior capabilities of neural networks in handling large-scale data and optimization problems, PINNs have garnered considerable attention since their introduction and have become important tools for solving PDEs. These methods utilize automatic differentiation to compute the penalty of an equation at a set of collocation points within the domain of interest, using it as an objective loss function that is optimized through backpropagation, thereby approximating the solution of the PDE. Compared to classical PDE numerical methods such as the finite element method and finite difference method, one significant advantage of PINNs is that they are mesh-free, making them easier to implement and particularly well-suited for high-dimensional problems and irregular domains. Currently, PINNs have been applied to solve a range of PDE problems, demonstrating their strengths effectively, such as inverse problems in fluid dynamics, optics and metamaterials [2, 3], or some other forward problems [4, 5, 6].
Despite the notable advantages demonstrated by PINNs in PDEs, there remains substantial room for improvement in several aspects. While PINNs have exhibited high-precision solutions across a variety of scenarios, their applicability is still limited. Notably, significant errors can arise when addressing equations such as the Allen-Cahn equation or the Burgers’ equation. Enhancements to the PINNs methodology can be pursued from multiple aspects. For instance, since the total loss function in PINNs is a weighted sum of errors from both the domain and boundary conditions, adjusting the weights of these terms may improve performance across different problems [7, 8, 9, 10]. In cases involving extensive solution domains, decomposition techniques can be employed to localize the training process, thereby enhancing optimization outcomes [11, 12, 13]. Regarding the choice of loss functions, although Mean Squared Error (MSE) is commonly used, alternative forms of loss functions may give better training results in certain circumstances [14, 10, 15]. Furthermore, strategies such as batch training and the design of specialized network architectures to deal with Dirichlet or periodic boundary conditions offer a promising direction to enhance the performance of PINNs [8, 16, 17]. In this paper, we focus on the limitations of PINNs in handling solutions with sharp local variations. This challenge, arising from singularities or high-gradient regions, is typically addressed by introducing adaptive sampling strategies for PINNs. However, existing sampling methods primarily only rely on the values of the loss function at local collocation points, without incorporating a rigorous a-posteriori error estimator. To address this issue, we propose the recovery-type a-posteriori estimator-enhanced adaptive PINN (R-PINN), which integrates the recovery-type a-posteriori estimator from classical numerical methods into PINNs. Furthermore, we introduce an adaptive sampling method, the Recovery-type estimator adaptive distribution (R-AD), to systematically implement adaptive refinement.
The sampling methods for collocation points in PINNs are generally categorized into uniform and non-uniform sampling. In typical uniform sampling, points are either randomly distributed or placed at equispaced intervals, remaining fixed throughout the training process. However, this static sampling strategy has notable limitations, as it fails to capture the model’s fine local features effectively, which can lead to convergence issues in complex problems settings. To address these limitations, Lu et al. first proposed an adaptive non-uniform sampling method [18]. Since then, various adaptive sampling methods for collocation points have been developed [19, 20, 21, 22, 23, 24, 20]. Among these, the Failure-informed PINN (FI-PINN) by Gao et al. [20] is particularly notable, utilizing a truncated Gaussian model for adaptive sampling. Another significant contribution is by Wu et al. [19], who reviewed and compared a range of sampling techniques. Their work covers both uniform sampling methods, such as equispaced, random, and low-discrepancy sequences, and non-uniform methods, including residual-based adaptive refinement (RAR) and residual-based adaptive distribution (RAD), offering valuable guidance on selecting appropriate sampling strategies.
While there are various adaptive PINN methods, they generally share a common approach: sampling the uniform points as background collocation points and then adding non-uniform points as the adaptive collocation points based on an a-posteriori error estimator. Most of these methods utilize the residual-type a-posteriori estimator, which calculates the error as the loss of the neural network function. For adaptive finite element methods, two primary types of estimator are widely recognized: residual-type [25, 26] and recovery-type [27, 28, 29], with the latter based on gradient recovery. Although extensive research has been conducted on adaptive PINNs, there has been limited work on gradient recovery estimators. Therefore, we introduce an adaptive PINN enhanced by the recovery-type a-posteriori estimator.
The rest of this paper is organized as follows. Section 2 provides preliminary background, introducing the fundamental algorithm of Physics-informed neural networks (PINNs) and existing sampling strategies for both uniform and non-uniform purposes. In Section 3, we present the detailed algorithm of the proposed Recovery-type estimator-enhanced PINN (R-PINN). Section 4 presents our results using R-PINN on four representative singular problems, illustrating its performance and providing a comparative analysis. Finally, Section 5 summarizes our conclusions and outlines potential directions for future research.
2 Preliminaries
In this section, we will present the requisite preliminary knowledge. We will begin with an overview of PINNs algorithm in solving both forward and inverse PDEs problems. Following this, we will discuss some strategies for collocation points sampling employed in the implementation of PINNs, including both uniform and non-uniform sampling methodologies.
2.1 PINNs for solving forward and inverse PDEs
Analogous to the application of basis functions for approximating true solutions in traditional numerical methods for PDEs, the Physics-informed neural networks (PINNs) algorithm utilizes deep neural networks to approximate the exact solutions of equations, primarily employing fully connected feed-forward neural networks. We denote a fully connected feed-forward neural network with $L$ layers as $\mathcal{N}^{L}:\mathbb{R}^{N_{0}}\rightarrow\mathbb{R}^{N_{k}}$, where the $k$-th layer comprises $N_{k}$ neurons. The weight matrix and bias vector of the $k$-th layer (for $1\leq k\leq L$) can be represented as $\bm{W}^{k}\in\mathbb{R}^{N_{k}\times N_{k-1}}$ and $\bm{b}^{k}\in\mathbb{R}^{N_{k}}$, respectively. The input vector is denoted as $\bm{x}\in\mathbb{R}^{N_{0}}$, and the output vector of the $k$-th layer is expressed as $\mathcal{N}^{k}(\bm{x})$, especially, $\mathcal{N}^{0}(\bm{x})=\bm{x}$. The activation function $\Phi$, applied during the feed-forward transition from the $(k-1)$-th layer to the $k$-th layer (for $1\leq k\leq L$), introduces non-linearity into the network, thereby enhancing its learning capacity. Consequently, a feed-forward neural network can be mathematically defined as
$\mathcal{N}^{k}(\bm{x})=\Phi(\bm{W}^{k}\mathcal{N}^{k-1}(\bm{x})+\bm{b}^{k}),\quad 1\leq k\leq L.$ (2.1)
Letting $\tilde{\bm{\Theta}}=\{\bm{W}^{k},\bm{b}^{k}\}\in\mathcal{V}$ be all parameters of the neural network, including weights and biases, where $\mathcal{V}$ denotes the parameter space, we can finally express the neural network as
$\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}}),$
where $\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ emphasizes the dependence of the output layer $\mathcal{N}^{L}(\bm{x})$ on $\tilde{\bm{\Theta}}$. Initial parameters are typically generated according to a specified probability distribution.
In the context of solving PDEs with PINNs, we consider a PDE defined over the domain $\Omega\in\mathbb{R}^{d}$ with parameters $\bm{\lambda}$,
$f(\bm{x};u(\bm{x}))=f\left(\bm{x};\frac{\partial u}{\partial x_{1}},\cdots,\frac{\partial u}{\partial x_{d}};\frac{\partial^{2}u}{\partial x_{1}^{2}},\cdots,\frac{\partial^{2}u}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)=0,\quad\bm{x}=(x_{1},x_{2},\cdots,x_{d})\in\Omega,$ (2.2)
subject to the boundary condition:
$\mathcal{B}(u,\bm{x})=0\quad\text{on}\ \partial\Omega,$
where $u(\bm{x})$ represents the value of the true solution at point $\bm{x}$.
In solving the forward problem, where the parameter $\bm{\lambda}$ in the equation is known, PINNs approximate the true solution $u$ using a neural network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})$ [1]. During this process, the neural network is iteratively trained to update its parameters $\tilde{\bm{\Theta}}$ in order to minimize the loss function, defined as,
$\mathcal{L}(\mathcal{T};\tilde{\bm{\Theta}})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}}),$ (2.3)
where
$\begin{split}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}})=\frac{1}{\lvert\mathcal{T}_{f}\rvert}\sum_{\bm{x}\in\mathcal{T}_{f}}\bigg{|}&f\left(\bm{x};\frac{\partial\hat{u}}{\partial x_{1}},\cdots,\frac{\partial\hat{u}}{\partial x_{d}};\frac{\partial^{2}\hat{u}}{\partial x_{1}\partial x_{1}},\cdots,\frac{\partial^{2}\hat{u}}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)\bigg{|}^{2},\\ &\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}})=\frac{1}{\lvert\mathcal{T}_{b}\rvert}\sum_{\bm{x}\in\mathcal{T}_{b}}\lvert\mathcal{B}(\hat{u},\bm{x})\rvert^{2},\end{split}$ (2.4)
and $w_{f}$ and $w_{b}$ are weighting factors. Here, $\mathcal{T}_{f}$ and $\mathcal{T}_{b}$ denote the sets of sampled points within the domain and on the boundary, respectively, which are commonly referred to as ”collocation points”. The combined set of collocation points is given by $\mathcal{T}=\mathcal{T}_{f}\cup\mathcal{T}_{b}$.
For certain problems, while the parameters $\bm{\lambda}$ in the governing equations remain unknown, the solutions are observable, thereby giving rise to inverse problems. To solve the inverse problem with PINNs, the loss function not only accounts for the deviations dictated by physical laws but also incorporates an additional term to capture the residual between observed solution $u(\bm{x})$ and predicted solution $\hat{u}(\bm{x})$ over a set of collocation points $\mathcal{T}_{i}$:
$\mathcal{L}_{i}(\mathcal{T}_{i};\tilde{\bm{\Theta}},\bm{\lambda})=\frac{1}{|\mathcal{T}_{i}|}\sum_{\bm{x}\in\mathcal{T}_{i}}|\hat{u}(\bm{x})-u(\bm{x})|^{2}.$ (2.5)
As a result, the total loss function for the inverse problem can be expressed as:
$\mathcal{L}(\mathcal{T};\tilde{\bm{\Theta}},\bm{\lambda})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}},\bm{\lambda})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}},\bm{\lambda})+w_{i}\mathcal{L}_{i}(\mathcal{T}_{i};\tilde{\bm{\Theta}},\bm{\lambda}),$ (2.6)
with the weight of the additional term $w_{i}$. Then the unknown parameters $\bm{\lambda}$ can be trained with the neural network parameters $\tilde{\bm{\Theta}}$ by the optimization process.
Note that in equations (2.3) and (2.6), it is crucial to position the set of collocation points within the domain of interest. While simple sampling methods, such as uniform and random sampling, are available, they are generally less effective in capturing the solution’s local features. Consequently, in certain cases, employing more advanced non-uniform sampling techniques becomes particularly important.
2.2 Sampling methods for PINNs
Typically, the sampling strategy in PINNs involves selecting a set of collocation points by uniform sampling prior to training the neural network, with the set remaining fixed throughout the training process. Commonly employed uniform sampling methods include equispaced uniform sampling and uniformly random sampling, while other uniform sampling techniques are also available [1, 4, 30]. In Wu. et al.’s work [19], six uniform sampling methods were systematically reviewed, including equispaced uniform grid (Grid), uniformly random sampling (Random), Latin hypercube sampling (LHS) [31, 32], Halton sequence (Halton) [33], Hammersley sequence (Hammersley) [34], and Sobol sequence (Sobol) [35]. Their comparative experiments demonstrated that the Sobol method generally provides superior performance. The Sobel sequence uses a base of two to form successively finer uniform partitions of the unit interval and then reorder the coordinates in each dimension. We show the difference over three different methods in Fig. 1.
Fig. 1: $900$ uniformly collocation points generated on $[0,1]^{2}$ with three different strategies.
In many cases, accuracy of numerical solution can be enough through uniform sampling strategies. However, for certain numerical problems characterized by large local gradients and highly distinctive solution features, such methods may fail to adequately capture local variations. Several adaptive methods have been proposed to overcome these limitations in recent years, primarily through modifications to the sampling strategies for the collocation points in PINNs. These adaptive sampling approaches can be broadly categorized into two types [19]: the Residual-based adaptive refinement (RAR) [18] and the Residual-based adaptive distribution (RAD) [21]. RAR targets regions with large residuals, where new adaptive collocation points are introduced at regular intervals during the training process to reduce local errors (see Algorithm 1).
Algorithm 1 Residual-based adaptive refinement (RAR)
1:  Sample background collocation points $\mathcal{T}$.
2:  Train the neural network with $\mathcal{T}$ for a certain number of epoches.
3:  while not reach the end do
4:     Sample a set of adaptive collocation points $\mathcal{S}_{0}$.
5:     Compute the PDE residuals of adaptive collocation points $\mathcal{S}_{0}$.
6:     $S\leftarrow m$ points with largest residuals from $\mathcal{S}_{0}$.
7:     $\mathcal{T}\leftarrow\mathcal{T}\cup\mathcal{S}$.
8:     Train the neural network for some epoches.
9:  end while
10:  Output the result.
While RAR only focuses on the area with large residuals and ignore the other place, RAD utilizes the probabilty density function $p(\bm{x})$ that has strong relationship with residual to resample all the collocation points, which can take the area both with large and small residual into care (see Algorithm 2). For example, we can take $p(\bm{x})$ proportional to the PDE residual $\epsilon(\bm{x})=|f(\bm{x};\hat{u}(\bm{x})|$,
[…]
More forms of $p(x)$ can be used here, such as it in paper[19],
$p(\bm{x})\propto\frac{\epsilon^{k}(\bm{x})}{\mathbb{E}[\epsilon^{k}(\bm{x})]}+c,$ (2.8)
where $k\geq 0,c\geq 0$ are two hyperparameters.
Algorithm 2 Residual-based adaptive distribution (RAD)
1:  Sample background collocation points $\mathcal{T}$.
2:  Train the neural network with $\mathcal{T}$ for a certain number of epoches.
3:  while not reach the end do
4:     $\mathcal{T}\leftarrow$ resampled collocation points according to the probability density function $p(\bm{x})$.
5:     Train the neural network for some epoches.
6:  end while
7:  Output the result.
3 Recovery-type estimator enhanced PINN
In this section, we present an adaptive PINNs algorithm developed by our team, which utilizes the recovery-type a-posteriori error estimator [27, 28, 29], a widely used technique in adaptive finite element methods. The concept of adaptive finite element method, first introduced by I. Babuška et al. in 1978 [25], has since seen significant advancements and remains integral to practical scientific computations. This approach extends classical finite element methods by incorporating a-posteriori error estimation and automated mesh refinement to iteratively improve the accuracy of approximate solutions, which can achieve the target accuracy with reduced computational effort, offering distinct advantages over traditional finite element methods, particularly when addressing problems involving singular solutions.
The fundamental concept of the adaptive finite element method lies in the use of a-posteriori error estimator to assess the numerical solution error within each mesh element, thereby guiding the mesh refinement and coarsening process. This adaptive mesh strategy shares a notable similarity with adaptive PINNs, where the distribution of collocation points is adjusted based on the error distribution. In adaptive finite element methods, commonly employed a-posteriori error estimators can be categorized into two main types: the first is based on residual type [25, 36], where the error is estimated by computing the discrepancy between the solution and the residual of the governing equation; the second type, recovery type, proposed by the Zienkiewicz and Zhu [27, 28, 29], is a gradient reconstruction-based method that provides more accurate error estimates by reconstructing the solution gradients.
In existing adaptive PINNs algorithms, the equation operator is typically used as the default a-posteriori error estimator, which corresponds to the residual-based error estimator in adaptive finite element methods. Our proposed algorithm seeks to introduce the recovery-type error estimator into the adaptive PINNs framework, with the aim of enriching and extending the theoretical foundation of adaptive PINNs, ultimately providing a more comprehensive approach to error estimation and control.
3.1 Recovery-type a-posteriori estimator
In 1987, O.C. Zienkiewicz and J.Z. Zhu introduced a pioneering a-posteriori error estimator based on gradient recovery for finite element theory, commonly referred to as the Z-Z method [27] or the recovery-type method. This approach, notable for its computational simplicity, intuitive understanding, and straightforward implementation within finite element software, quickly gained widespread acceptance within the engineering community. Its rapid adoption represented a significant milestone, facilitating the transition from theoretical research to practical application in the domain of finite element a-posteriori error estimation.
In the context of a specific recovery-type a-posteriori error estimation framework, a typical problem in the field of elasticity is to be considered as an example, which is encapsulated within the following system of equations
$\bm{L}\bm{u}=\bm{S}^{T}\bm{DS}\bm{u}=\bm{f}\quad\text{in}\ \Omega,$ (3.1)
which is subject to suitable boundary conditions on the domain’s boundary $\partial\Omega$. Here, $\bm{u}$ represents the displacement vector, $\bm{L}$ is the differential operator intrinsic to the theory of elasticity, and the matrix operator $\bm{S}$ defines the strain tensor $\bm{\epsilon}$ by
$\bm{\epsilon}=\bm{S}\bm{u}.$ (3.2)
Furthermore, the elasticity matrix $\bm{D}$ is utilized to determine the stress tensor $\bm{\sigma}$, as delineated by the equation:
$\bm{\sigma}=\bm{D}\bm{\epsilon}.$ (3.3)
The finite element solution to the aforementioned elasticity problem is expressed in the following form
$\bm{u}_{h}=\bm{N}\bar{\bm{u}}_{h},$ (3.4)
where $\bar{\bm{u}}_{h}$ denotes the vector of nodal displacements, and $\bm{N}$ represents the set of approximate polynomial shape functions that interpolate the displacement field within the finite element framework Subsequently, the stress tensor $\bm{\sigma}_{h}$, is computed using the approximate displacement field $\bm{u}_{h}$ with
$\bm{\sigma}_{h}=\bm{DSu}_{h}.$ (3.5)
The recovery-type a-posteriori error estimator is characterized by focusing on estimating the numerical error by approximating the gradient recovery of the numerical solution $\bm{\sigma}^{*}$. Therefore, the gradient recovery technique plays a key role for the estimator. With the development of recovery-type a-posteriori error estimators, there have been many gradient recovery methods, the typical ones are the transformation method and nodal averaging method.
Fig. 2: The finite element method solution and the smoothened solution of stress for a one-dimensional problem.
  • 1.
    The transformation method
    For this method, it is assumed that the stress
    $\bm{\sigma}^{*}=\bm{N}\bar{\bm{\sigma}}^{*},$ (3.6)
    and the method seeks to minimize the discrepancy between the smooth stress $\bm{\sigma}^{*}$ and the stress of the finite element $\bm{\sigma}_{h}$ by enforcing the condition
[…]
    By combining equations (3.6) and (3.7), we can derive the following expression for $\bar{\bm{\sigma}}^{*}$
[…]
    This method is highly accurate but computationally expensive due to the need to solve the additional system of equations.
  • 2.
    The nodal averaging method
    In this simpler approach, the smooth stress $\bar{\sigma}^{*}$ is obtained by averaging the finite element stresses $\bm{\sigma}_{h,i}$ at the nodes
    $\bm{\sigma}^{*}=\frac{\left(\sum_{i=1}^{M}\bm{\sigma}_{h,i}\right)}{M}$ (3.9)
    where $M$ is the number of elements connected with the node $i$.
    This method is straightforward and computationally efficient, although it may not achieve the same level of accuracy as the transformation method. Nevertheless, it performs well in practical applications.
As for error estimation, two types of errors are usually considered. One is the error in $\bm{\sigma}$ which can be expressed as
$\lVert\bm{e}_{\sigma}\rVert=\lVert\bm{\sigma}-\bm{\sigma}_{h}\rVert.$ (3.10)
The other is the recovery-type estimator of the error is similarly defined in the appropriate norm as
$\lVert\tilde{\bm{e}}_{\sigma}\rVert=\lVert\bm{\sigma}^{*}-\bm{\sigma}_{h}\rVert,$ (3.11)
and its efficiency is gauged by the efficacy index defined as
$\theta=\frac{\lVert\tilde{\bm{e}}_{\sigma}\rVert}{\lVert\bm{e}_{\sigma}\rVert}.$ (3.12)
The confines of this efficacy index are elucidated through the subsequent theorem (as detailed in [29]):
$1-\frac{\lVert\bm{e}_{\sigma}^{*}\rVert}{\lVert\bm{e}_{\sigma}\rVert}\leq\theta\leq 1+\frac{\lVert\bm{e}_{\sigma}^{*}\rVert}{\lVert\bm{e}_{\sigma}\rVert},$ (3.13)
with the error in the reconstructed solution characterized as
$\lVert\bm{e}_{\sigma}^{*}\rVert=\lVert\bm{\sigma}-\bm{\sigma}^{*}\rVert.$ (3.14)
It is evident that the efficacy index approaches unity when the precision of the reconstructed solution significantly outperforms that of the original finite element solution.
Should the reconstructed solution exhibit superconvergence, that is,
$\lVert\bm{e}_{\sigma}^{*}\rVert=C^{*}h^{p+\alpha},$ (3.15)
where $\alpha$ is a positive constant and $h$ represents the element size, and
$\lVert\bm{e}_{\sigma}\rVert=Ch^{p},$ (3.16)
then the bounds for the efficacy index are
$1-\tilde{C}h^{\alpha}\leq\theta\leq 1+\tilde{C}h^{\alpha},$ (3.17)
with $\tilde{C}=C^{*}/C$, rendering the error estimator asymptotically precise. This estimation underscores the efficiency of the recovery-type a-posteriori estimator. Given that the error estimation revolves around stress $\bm{\sigma}$, which is contingent upon the gradients of the solution $\bm{u}$, this class of estimator may also be regarded as a form of gradient recovery.
3.2 Algorithm of Recovery-type estimator enhanced PINN
In this section, we introduce the algorithm of recovery-type estimator enhanced adaptive PINN (R-PINN). The success of the adaptive PINNs algorithm hinges on two crucial components: the posterior error estimation method and the adaptive sampling strategy. The posterior error estimation predicts the discrepancy between the numerical and true solutions, while the adaptive sampling strategy adjusts the distribution of collocation points based on this error, allowing for more focused and efficient training in areas that need it most. Section 3.1 provides a comprehensive explanation of the posterior error estimation method. Therefore, for our R-PINN, the key is how to combine PINNs with this kind of posterior error estimator to realize the adaptive process.
To solve the equation (2.2), we utilize a shape-regular triangulation $\mathcal{T}_{h}$ formed by a family of partitions of the domain $\Omega$, which allows us to effectively implement the recovery-type a-posteriori estimator. We define the operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ as the recovery-type estimator used in our algorithm, employing a nodal averaging method for gradient recovery (other gradient recovery methods can also be utilized). The approximated solution $\hat{u}_{\tilde{\Theta}}$ predicts the solution values at all nodes within the mesh $\mathcal{T}_{h}$, which then facilitates gradient recovery and error computation for each element $K\in\mathcal{T}_{h}$ through the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$.
Fig. 3: The flow chart of R-PINN.
Another crucial aspect is the selection of collocation points and the adaptive adjustment of their distribution. Similar to many adaptive PINNs algorithms, we start with $N_{1}$ background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ for pre-training. During the adaptive training phase, we introduce $N_{2}$ adaptive collocation points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ alongside the background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$, updating the adaptive collocation points after each training batch. For background collocation points, we utilize a Sobol sequence as described in section 2.2. For the adaptive collocation points, we propose a novel approach called the recovery-type estimator based adaptive distribution (RE-AD).
Unlike common sampling methods, the RE-AD technique is element-based, as the computed a-posteriori errors are distributed across elements. This means we only need to specify the number of adaptive collocation points to sample in each element, allowing for random sampling within each element. Therefore, it’s vital to allocate adaptive collocation points intelligently to target areas with significant errors without creating redundancy. To achieve this, we design an iterative process: we define a tolerance $\epsilon$ such that $\frac{N_{2}}{N_{e}}\leq\epsilon<1$, where $N_{e}$ is the total number of elements. Initially, we select $L_{1}=[\epsilon N_{e}]$ elements that exhibit the largest errors in $\mathcal{T}_{h}$, which we refer to as the sequence $\mathcal{S}_{1}$. Adaptive collocation points are sampled from $\mathcal{S}_{1}$ in proportion to the errors. For instance, for the $i$-th element in $\mathcal{S}_{1}$, the number of adaptive collocation points assigned is $[N_{2}\frac{e_{i}}{\sum_{j}e_{j}}]$. Since this process rounds down, some adaptive collocation points may remain unallocated. We then repeat this process by selecting $L_{2}=[\epsilon L_{1}]$ of the largest elements and allocating adaptive collocation points according to their errors, continuing until no points are redundant. The RE-AD sampling method differs from RAR but is more akin to RAD, as it maintains a fixed total number of adaptive collocation points while only adjusting their distribution during each iteration.
Algorithm 3 Recovery-type estimator based adaptive distribution (RE-AD)
1:  Set the number of redundant adaptive collocation points $N_{r}=N_{2}$.
2:  Denote $L_{0}=N_{e}$.
3:  $k\leftarrow 1$.
4:  while $N_{r}$ not 00 do
5:     Select $L_{k}=[\epsilon L_{k-1}]$ elements with the largest errors as $\mathcal{S}_{k}$.
6:     Designate adaptive collocation points to $\mathcal{S}_{k}$.
7:     Update the number of redundant adaptive collocation points $N_{r}$.
8:     $k\leftarrow k+1$.
9:  end while
10:  Output adaptive collocation points $\{\bm{x}_{a}\}_{i=1}^{N_{2}}$.
To introduce the specific procedures of the recovery-type estimator enhanced adaptive PINN (R-PINN), we utilize the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ along with the recovery-type estimator based adaptive distribution (RE-AD) sampling strategy. Initially, we use a set of background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ to train the neural network, aiming to obtain a rough approximation of the true solution, which serves as the pre-training step. Following the pre-training phase, an adaptive training iteration process begins. In each iteration, the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ is used to compute the errors of the approximated solution $\hat{u}_{\tilde{\Theta}}(\bm{x})$ on the mesh $\mathcal{T}_{h}$. Based on these errors, the RE-AD sampling method is then employed to strategically distribute additional adaptive collocation points. These adaptive collocation points are combined with the background collocation points to retrain the neural network. This iterative process continues until the accuracy improvements become minimal, indicating that the network has converged to a satisfactory approximation of the solution. Since the number of collocation points can always stay constant, the algorithm satisfies the characteristics of self-adaptation while ensuring accurate solution.
Algorithm 4 Recovery-type estimator enhanced PINN (R-PINN)
0:  The number of adaptive samples and background samples $N_{a}$ and $N_{b}$, the maximum steps number $M$
1:  Generate $N_{1}$ background samples $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ with Sobol sequence.
2:  Pre-train the network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ with background samples $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$.
3:  Set the number of iteration $k\leftarrow 1$ and give the max number of iterations $M$.
4:  while $k\leq M$ do
5:     Generate $N_{2}$ adaptive samples $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ with RE-AD method.
6:     Retrain the network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ with the dataset $\mathcal{D}=\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}\cup\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$.
7:  end while
[…]
4 Numerical experiments
In this section, we present numerical experiments demonstrating the effectiveness of our proposed algorithm. Our evaluation compares three different kinds of PINNs: the Recovery-type estimator enhanced PINN (R-PINN), Failure-informed PINN (FI-PINN) [20], and standard PINN as a baseline.
4.1 Experimental setup
To ensure a fair comparison, we maintain identical parameters and training processes across all methods for the same number of iterations, varying only the number and distribution of adaptive collocation points. The recovery-type a-posteriori estimation is implemented using FEALPy [37].
All experiments employ a fully-connected neural network with $7$ hidden layers, each containing 20 neurons with tanh for the activation function unless otherwise specified. The training process comprises two phases: an initial pre-training phase utilizing the L-BFGS optimizer with a learning rate of $0.1$, conducted for a problem-dependent number of epochs, followed by an adaptive training phase consisting of $M$ adaptive iterations. Each iteration is constrained to $5000$ epochs while employing the same optimization method.
For collocation points, we use $N_{1}=2000$ background collocation points and $N_{2}=100$ adaptive collocation points for R-PINN unless specified, while FI-PINN requires $N_{2}=500$ adaptive collocation points to achieve convergence unless specified. For the standard PINN, only $N_{1}=2000$ background collocation points are used during training, without incorporating additional adaptive collocation points. However, the training process remains the same as that of the two adaptive PINNs, with each adaptive iteration also consisting of 5000 epochs. All methods utilize $200$ boundary collocation points. Notably, our approach achieves comparable or superior accuracy despite using fewer total epochs and adaptive collocation points than related studies. The larger $N_{2}$ value for FI-PINN is necessary as it fails to achieve adaptivity with fewer points. Furthermore, while our method maintains a fixed number of adaptive collocation points, FI-PINN accumulates points throughout the adaptive iteration process. To evaluate the accuracy of our method, we employ two error metrics: the relative $L^{2}$ error,
$\frac{\lVert\hat{u}{\tilde{\Theta}}-u\rVert_{2}}{\lVert u\rVert_{2}},$ (4.1)
and the $L^{\infty}$ error, defined as $\lVert\hat{u}{\tilde{\Theta}}-u\rVert_{\infty}$. Here, $\hat{u}_{\tilde{\Theta}}$ denotes the numerical solution obtained by our method, and $u$ represents the exact solution.
4.2 Poisson equation
Consider the following Poisson equation:
$\begin{split}-\Delta u(x,y)=&f(x,y),\quad(x,y)\ \text{in}\ \Omega,\\ u(x,y)=&g(x,y),\quad(x,y)\ \text{on}\ \partial\Omega,\end{split}$ (4.2)
where the domain $\Omega$ is $=[0,1]\times[0,1]$, and the true solution is
$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}.$ (4.3)
The solution exhibits a sharp peak near the point $(0.5,0.5)$, which is challenging for standard PINN to capture accurately. In this experiment, we pre-train the network for $10000$ iterations and set the maximum number of adaptive iterations to $M=5$.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $3.386$ $0.0990$ $100$ $2.717$ $0.6541$ $500$ $0.809$ $0.6270$
$2$ $0.732$ $0.0282$ $100$ $1.514$ $0.4593$ $1000$ $0.689$ $0.6180$
$3$ $0.125$ $0.0260$ $100$ $1.157$ $0.3095$ $1500$ $0.650$ $0.5683$
$4$ $0.285$ $0.0088$ $100$ $0.294$ $0.0686$ $2000$ $0.879$ $0.5689$
$5$ $0.139$ $0.0086$ $100$ $0.068$ $0.0335$ $2500$ $0.682$ $0.5570$
Table 1: The relative $L^{2}$ error and $L^{\infty}$ error by two methods solving Poisson equation.
R-PINN
(a) *
FI-PINN
(b) *
Fig. 4: The distribution of adaptive collocation points in the first $2$ adaptive iterations with two adaptive PINN methods.
Fig. 4 illustrates the distribution of adaptive collocation points by two adaptive methods in the first two adaptive iterations. It is evident that all adaptive collocation points of our model converge to the neighborhood of the point $(0.5,0.5)$, while adaptive collocation points of FI-PINN are more dispersed. This shows the strong effectiveness of our algorithm in handling problems with significant local variations. Table 1 compares the relative $L^{2}$ error and $L^{\infty}$ error among R-PINN, FI-PINN and standard PINN during the adaptive iterations where $N_{\text{iter}}$ represents the number of adaptive iterations. The relative $L^{2}$ error for R-PINN reaches $0.139$, and the $L^{\infty}$ error reaches $0.0086$. In contrast, the errors obtained by FI-PINN decrease slowly at the beginning, but drop fast later, reaching the relative $L^{2}$ error of $0.068$ and $L^{\infty}$ error of $0.0335$. It can be observed that R-PINN can converge faster than FI-PINN. While, R-PINN can reach a smaller relative $L^{2}$ error in this case with the exponentially accumulated adaptive collocation points, which is one of its advantage. As for the normal PINN, the errors obtained show little to no decrease, indicating that the normal PINN fails to converge effectively. Similar results are observed in Fig. 6, where the lowest relative $L^{2}$ error for our algorithm is below $1\times 10^{-1}$. In the left panel of Fig. 5, we show the absolute errors throughout the entire domain, clearly demonstrating that the errors are much smaller with R-PINN. The middle panel of Fig. 5 illustrates the cross-section at $y=0.5$. The predicted curve using R-PINN fits the true solution very well, while the solution by FI-PINN has a little obvious error at the foot of the peak. The final solutions with three methods are plotted in the right of Fig. 5.
R-PINN
(a) *
FI-PINN
(b) *
standard PINN
(c) *
Fig. 5: The prediction err (left), the cross section at $y=0.5$ (middle) and the predicted solution of three methods solving Poisson equation.
Fig. 6: The relative $L^{2}$ errors of three methods in the optimization process of solving Poison equation.
4.3 Burgers’ equation
We consider the following Burgers’ equation:
$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\frac{0.01}{\pi}\frac{\partial^{2}u}{\partial x^{2}},\\ &u(x,0)=-\sin(\pi x),\\ &u(t,-1)=u(t,1)=0\end{split}$ (4.4)
where the domain is $(x,t)\in[-1,1]\times[0,1]$. The solution to this Burgers’ equation exhibits a sharp mutation along $x=0$, making it prone to errors and offsets. For this test, we pre-train the network for $5000$ iterations and use $M=4$ steps of adaptive processes.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.0086$ $0.0735$ $100$ $0.2278$ $1.8142$ $200$ $0.0136$ $0.1583$
$2$ $0.0107$ $0.0964$ $100$ $0.0131$ $0.1263$ $400$ $0.0144$ $0.1685$
$3$ $0.0071$ $0.0673$ $100$ $0.0064$ $0.0576$ $600$ $0.0174$ $0.2099$
$4$ $0.0059$ $0.0439$ $100$ $0.0089$ $0.0846$ $800$ $0.0197$ $0.2418$
Table 2: The relative $L^{2}$ error and $L^{\infty}$ error by three methods solving Burgers’ equation.
R-PINN
(a) *
FI-PINN
(b) *
Fig. 7: The distribution of adaptive collocation points in the first $2$ adaptive iterations with two adaptive PINN methods solving Burgers’ equation.
The adaptive collocation points generated during the first two adaptive training processes by two adaptive PINN method are illustrated in Fig. 7. We observe that the adaptive collocation points successfully cluster around $x=0$ with both methods, where the error is largest (as seen in Fig. 8). This clustering significantly improves the local error. The exact relative $L^{2}$ error and $L^{\infty}$ error are listed in Table 2. As shown, the final relative $L^{2}$ error and $L^{\infty}$ error for R-PINN are $0.0059$ and $0.0439$, respectively, while those for FI-PINN are $0.0089$ and $0.0846$. The prediction err of standard PINN keeps at the save level during the training, where relative $L^{2}$ error and $L^{\infty}$ reach $0.0197$ and $0.2418$ in the final. Similar to the experiment of Poisson equation, it takes a long iterations for FI-PINN before the error jumps quickly, while R-PINN can reduce the errors at once. The errors obtained by R-PINN and FI-PINN are one order of magnitude lower than those obtained by the standard PINN. Such behaviour can also be observed in the Fig. 10, where the relative $L^{2}$ errors are illustrated specifically, showing that the lowest error can reach below $1\times 10^{-2}$. We also show the absolute errors across the entire domain and the cross-section at $t=1$ in Fig. 8. The predicted solutions of two adaptive methods are illustrated in Fig. 9.
R-PINN
(a) *
FI-PINN
(b) *
standard PINN
(c) *
Fig. 8: The prediction err (left) and the cross section at $t=1$ (right) of three methods solving Burgers’ equation.
Fig. 9: The predicted solution (R-PINN at the left and FI-PINN at the right).
Fig. 10: The relative $L^{2}$ errors of three methods in the optimization process of solving Burgers’ equation.
4.4 Two-dimensional problem with two peaks
Considering the following two-dimensional problem
$\begin{split}-\nabla\cdot[u(x,y)\nabla(x^{2}+y^{2})]+\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in\Omega,\\ u(x,y)=b(x,y),\quad(x,y)\in\partial\Omega,\end{split}$ (4.5)
where $\Omega=[-1,1]\times[-1,1]$. The true solution is
$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}+e^{-1000[(x+0.5)^{2}+(y+0.5)^{2}]},$ (4.6)
which has two peaks at $(-0.5,-0.5)$ and $(0.5,0.5)$. This test mainly can show the ability of R-PINN to deal with multiple peaks. For this experiment, we pre-train our network for the period of $10000$ iterations and set the maximum number of adaptive iterations as $M=4$. Since two peaks problem is challenging for neural network to approximate, we use $5$ hidden layers with $64$ neurons in each layer to enhance its capability. Moreover, we use $N_{2}=150$ adaptive collocation points for R-PINN, and $N_{2}=750$ added adaptive collocation points for FI-PINN.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.6778$ $0.1228$ $150$ $0.4372$ $0.4845$ $750$ $0.7912$ $0.6651$
$2$ $0.1202$ $0.0198$ $150$ $0.5220$ $0.3429$ $1500$ $0.8020$ $0.6475$
$3$ $0.1080$ $0.0259$ $150$ $0.4773$ $0.2233$ $2250$ $0.8012$ $0.6483$
$4$ $0.0840$ $0.0127$ $150$ $0.3969$ $0.2263$ $3000$ $0.8342$ $0.6625$
Table 3: The relative $L^{2}$ error and $L^{\infty}$ error by two methods solving two-dimensional problem with two peaks.
R-PINN
(a) *
FI-PINN
(b) *
Fig. 11: The distribution of adaptive collocation points in the first two adaptive iterations of two adaptive methods solving the two-dimensional problem with two peaks.
R-PINN
(a) *
FI-PINN
(b) *
standard PINN
(c) *
Fig. 12: The predicted err (left), the cross section at $y=-0.5$ (middle) and the cross section at $y=0.5$ (right) of three methods solving the two-dimensional problem with two peaks.
Fig. 13: The predicted solution (R-PINN at the left, FI-PINN at the middle and standard PINN at the right) of the two-dimensional problem with two peaks.
Fig. 14: The relative $L^{2}$ errors of three methods in the optimization process of solving the two-dimensional problem with two peaks.
We present the distribution of adaptive collocation points of two adaptive methods across the first two adaptive iterations in Fig. 11. It is evident that these points converge around the vicinities of $(0.5,0.5)$ and $(-0.5,-0.5)$ with R-PINN, demonstrating the capability of our algorithm to effectively handle problems characterized by multiple regions of abrupt changes. However, the distribution of adaptive collocation points with FI-PINN seems not so effective to deal with the two peaks. The reason of it might be a truncated Gaussian distribution is used, so that problems of more than one regions with large error are difficult to handle. This problem is also reflected in the training process. Table 3 compares the relative $L^{2}$ and $L^{\infty}$ errors among R-PINN, FI-PINN and standard PINN during the adaptive iterations. For R-PINN, the relative $L^{2}$ error reduces to $0.0840$, and the $L^{\infty}$ error decreases to $0.0127$. In contrast, the relative $L^{2}$ error associated with FI-PINN can only reach $0.3969$ and the $L^{\infty}$ can get to $0.2263$, which is just a little bit lower than those with standard PINN. Similar trends are observed in Fig. 14, where the lowest relative $L^{2}$ error for R-PINN falls below $1\times 10^{-1}$.
Fig. 12 provides additional insights. The left panel illustrates the absolute errors across the entire domain, revealing significantly smaller errors for R-PINN. The middle and right panels display cross-sectional views at $y=-0.5$ and $y=0.5$, respectively, showing that the predictions from R-PINN align closely with the exact solution, whereas there are large errors near two peaks with FI-PINN and the standard PINN can only approximate the features. Finally, the numerical solutions obtained by the three methods are visualized in Fig. 13.
4.5 Wave equation
For this test, we consider a wave equation as follow,
$\begin{split}&\frac{\partial^{2}u}{\partial t^{2}}-3\frac{\partial^{2}u}{\partial x^{2}}=0,\quad(t,x)\in[0,6]\times[-5,5],\\ &u(0,x)=\frac{1}{\cosh(2x)}-\frac{0.5}{\cosh(2(x-10))}-\frac{0.5}{\cosh(2(x+10))},\\ &\frac{\partial u}{\partial t}(0,x)=0,\\ &u(t,-5)=u(t,5)=0,\end{split}$ (4.7)
where the true solution is
$u(t,x)=\frac{0.5}{\cosh(2(x-\sqrt{3}t))}-\frac{0.5}{\cosh(2(x-10+\sqrt{3}t))}+\frac{0.5}{\cosh(2(x+\sqrt{3}t))}-\frac{0.5}{\cosh(2(x+10-\sqrt{3}t))}.$ (4.8)
Fig. 15: The plot of the exact solution.
R-PINN
(a) *
FI-PINN
(b) *
standard PINN
(c) *
Fig. 16: The predicted solution (left), predicted error (middle) and the distribution of adaptive collocation points (right) of three methods solving the wave equation.
Fig. 17: The relative $L^{2}$ errors of three methods in the optimization process of solving the Poison equation the wave equation.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $L^{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.0572$ $0.0761$ $100$ $0.1110$ $0.1590$ $500$ $0.3448$ $0.4554$
$2$ $0.0076$ $0.0098$ $100$ $0.0317$ $0.0491$ $1000$ $0.3158$ $0.4214$
Table 4: The relative $L^{2}$ error and $L^{\infty}$ error by three methods solving the wave equation.
In this test, the R-PINN method quickly reaches its performance limit. Therefore, we adopt a single adaptive iteration process ($M=2$) and pre-train the network for 5000 iterations. We show the relative $L^{2}$ errors and $L^{\infty}$ errors in the Table 4. The relative $L^{2}$ error and $L^{\infty}$ error can reach $0.0076$ and $0.0098$ with R-PINN, which are obviously lower than the other two methods. The detailed decreasing process is convenient to be seen in Fig. 17 The exact solution is depicted in Fig. 15. The left panel of Fig. 16 compares the predicted solutions of the three methods. The middle panel of Fig. 16 illustrates the absolute errors across the domain, showing that the errors associated with R-PINN are smaller than those with FI-PINN and standard PINN. Finally, the right panel of Fig. 16 displays the adaptive collocation points for two adaptive methods and random points for the standard PINN. From the right panel, we can see that the same problem occurs for FI-PINN when dealing with multiple large errors region, so it can’t handle well.
5 Conclusion
In this paper, we present recovery-type a-posteriori estimator enhanced adaptive PINN that combines the PINN with a recovery type estimator which is popularly used in the adaptive finite element method. Our innovation lies in the use of the recovery-type a-posteriori error estimator, which is different from the residual of PINNs, and the design of a tailored sampling method R-AD. We test the performance of R-PINN in several PDE problems with sharp or singular solutions, which normal PINN struggles to handle effectively. The results demonstrate that R-PINN achieves significant improvements over normal PINN. More studies can be done to analyze the performance of R-PINN for more complex problems and we shall focus on and analyze these issues in our future work.

### v2

@@
 In recent years, with the advancements in machine learning and neural networks, algorithms using physics-informed neural networks (PINNs) to solve PDEs have gained widespread applications.
-While these algorithms are well-suited for a wide range of equations, they often exhibit suboptimal performance when applied to equations with large local gradients, resulting in substantial localized errors.
+While these algorithms are well-suited for a wide range of equations, they often exhibit a suboptimal performance when applied to equations with large local gradients, resulting in substantially localized errors.
 To address this issue, this paper proposes an adaptive PINN algorithm designed to improve accuracy in such cases.
-The core idea of the algorithm is to adaptively adjust the distribution of collocation points based on the recovery-type a-posterior error of the current numerical solution, enabling a better approximation of the true solution.
+The core idea of the algorithm is to adaptively adjust the distribution of collocation points based on the recovery-type a-posteriori error of the current numerical solution, enabling a better approximation of the true solution.
 This approach is inspired by the adaptive finite element method.
-By combining the recovery-type a-posteriori estimator, a gradient-recovery estimator commonly used in the adaptive finite element method (FEM) with PINNs, we introduce the Recovery-type a-posteriori estimator enhanced adaptive PINN (R-PINN) and compare its performance with a typical adaptive PINN algorithm, FI-PINN.
-Our results demonstrate that R-PINN achieves faster convergence with fewer adaptive points and significantly outperforms in the cases with multiple regions of large errors than FI-PINN.
-Notably, our method is a hybrid numerical approach for solving partial differential equations, integrating adaptive FEM with PINNs.
+By combining the recovery-type a-posteriori estimator, a gradient-recovery estimator commonly used in the adaptive finite element method (FEM), with PINNs, we introduce the recovery-type a-posteriori estimator enhanced adaptive PINN (R-PINN) and compare its performance with a typical adaptive sampling PINN, failure-informed PINN (FI-PINN), and a typical adaptive weighting PINN, residual-based attention in PINN (RBA-PINN) as a baseline.
+Our results demonstrate that R-PINN achieves faster convergence with fewer adaptively distributed points and outperforms the other two PINNs in the cases with regions of large errors.
 keywords:
-Adaptive PINN , Recovery-type a-posteriori estimator , Machine learning , Adaptive FEM
+physics-informed neural networks , adaptive sampling , recovery-type estimator , gradient recovery
 1 Introduction
 The physics-informed neural networks (PINNs) algorithm is a data-driven approach for solving partial differential equations (PDEs) based on deep learning frameworks [26].
-Due to its distinct methodology, which diverges significantly from traditional PDE numerical techniques, and by leveraging the superior capabilities of neural networks in handling large-scale data and optimization problems, PINNs have garnered considerable attention since their introduction and have become important tools for solving PDEs.
-These methods utilize automatic differentiation to compute the penalty of an equation at a set of collocation points within the domain of interest, using it as an objective loss function that is optimized through backpropagation, thereby approximating the solution of the PDE.
+Due to its distinct methodology, which diverges significantly from traditional numerical techniques, and by leveraging the superior capabilities of neural networks in handling large-scale data and optimization problems, PINNs have garnered considerable attention ever since their introduction and have become important tools for solving PDEs.
+These methods utilize automatic differentiation to compute the residual of an equation at a set of collocation points within the domain of interest, using it as an objective loss function that is optimized through backpropagation, thereby approximating the solution of the PDE.
 Compared to classical PDE numerical methods such as the finite element method and finite difference method, one significant advantage of PINNs is that they are mesh-free, making them easier to implement and particularly well-suited for high-dimensional problems and irregular domains.
-Currently, PINNs have been applied to solve a range of PDE problems, demonstrating their strengths effectively, such as inverse problems in fluid dynamics, optics and metamaterials [2, 3], or some other forward problems [4, 5, 6].
-Despite the notable advantages demonstrated by PINNs in PDEs, there remains substantial room for improvement in several aspects.
+Currently, PINNs have been applied to solve a range of PDE problems, including fluid mechanics, which has been reviewed comprehensively by Cai et al.
+[5], as well as fields such as optics, metamaterials and others [27, 6, 24, 38, 12].
+Despite the notable advantages demonstrated by PINNs in PDEs, there remains a substantial room for improvement.
 While PINNs have exhibited high-precision solutions across a variety of scenarios, their applicability is still limited.
@@
 Enhancements to the PINNs methodology can be pursued from multiple aspects.
-For instance, since the total loss function in PINNs is a weighted sum of errors from both the domain and boundary conditions, adjusting the weights of these terms may improve performance across different problems [7, 8, 9, 10].
+For instance, since the total loss function in PINNs is a weighted sum of errors from both the domain and boundary conditions, adjusting the weights of these terms may improve performance across different problems [1, 32, 19, 33, 10].
+A representative example is the residual-based attention in PINN [1], which selects weighting multipliers for the loss function and has been shown to significantly enhance the performance of PINNs.
 In cases involving extensive solution domains, decomposition techniques can be employed to localize the training process, thereby enhancing optimization outcomes [22, 28, 16].
-Regarding the choice of loss functions, although Mean Squared Error (MSE) is commonly used, alternative forms of loss functions may give better training results in certain circumstances [14, 10, 15].
-Furthermore, strategies such as batch training and the design of specialized network architectures to deal with Dirichlet or periodic boundary conditions offer a promising direction to enhance the performance of PINNs [8, 16, 17].
+Regarding the choice of loss functions, although the mean squared error (MSE) is commonly used, alternative forms of loss functions may give better training results in certain circumstances [10, 25, 20].
+Furthermore, strategies, such as batch training and the design of specialized network architectures to deal with Dirichlet or periodic boundary conditions, offer a promising direction to enhance the performance of PINNs [19, 17, 7].
 In this paper, we focus on the limitations of PINNs in handling solutions with sharp local variations.
-This challenge, arising from singularities or high-gradient regions, is typically addressed by introducing adaptive sampling strategies for PINNs.
-However, existing sampling methods primarily only rely on the values of the loss function at local collocation points, without incorporating a rigorous a-posteriori error estimator.
-To address this issue, we propose the recovery-type a-posteriori estimator-enhanced adaptive PINN (R-PINN), which integrates the recovery-type a-posteriori estimator from classical numerical methods into PINNs.
-Furthermore, we introduce an adaptive sampling method, the Recovery-type estimator adaptive distribution (R-AD), to systematically implement adaptive refinement.
+This challenge, arising from singularities or high-gradient regions, is typically addressed by introducing adaptive sampling.
+However, existing sampling methods primarily rely on the PDE loss at local collocation points, which can be regarded as a residual-type estimator.
+Nevertheless, alternative efficient estimators remain unexplored for adaptive PINNs.
+Therefore, we propose the recovery-type a-posteriori estimator enhanced adaptive PINN (R-PINN), which integrates the recovery-type a-posteriori estimator from classical numerical methods into PINNs.
+Furthermore, we introduce an adaptive sampling method, the recovery-type estimator adaptive distribution (RecAD), which performs adaptive refinement based primarily on the recovery-type error estimator.
 The sampling methods for collocation points in PINNs are generally categorized into uniform and non-uniform sampling.
@@
 first proposed an adaptive non-uniform sampling method [18].
-Since then, various adaptive sampling methods for collocation points have been developed [19, 20, 21, 22, 23, 24, 20].
+Since then, various adaptive sampling methods for collocation points have been developed [35, 9, 23, 8, 37, 15].
 Among these, the failure-informed PINN (FI-PINN) by Gao et al.
@@
 Their work covers both uniform sampling methods, such as equispaced, random, and low-discrepancy sequences, and non-uniform methods, including residual-based adaptive refinement (RAR) and residual-based adaptive distribution (RAD), offering valuable guidance on selecting appropriate sampling strategies.
-While there are various adaptive PINN methods, they generally share a common approach: sampling the uniform points as background collocation points and then adding non-uniform points as the adaptive collocation points based on an a-posteriori error estimator.
+While there are various adaptive PINNs methods, they generally share a common feature: sampling the uniform points as background collocation points and then adding non-uniform points as the adaptively distributed points based on an error estimator.
 Most of these methods utilize the residual-type a-posteriori estimator, which calculates the error as the loss of the neural network function.
 For adaptive finite element methods, two primary types of estimator are widely recognized: residual-type [3, 2] and recovery-type [39, 40, 41], with the latter based on gradient recovery.
-Although extensive research has been conducted on adaptive PINNs, there has been limited work on gradient recovery estimators.
+Although extensive research has been conducted on residual-type adaptive PINNs [35, 9], the use of gradient recovery-based estimators remains unexplored.
 Therefore, we introduce an adaptive PINN enhanced by the recovery-type a-posteriori estimator.
@@
 Section 2 provides preliminary background, introducing the fundamental algorithm of physics-informed neural networks (PINNs) and existing sampling strategies for both uniform and non-uniform purposes.
-In Section 3, we present the detailed algorithm of the proposed Recovery-type estimator-enhanced PINN (R-PINN).
+Section 3 presents the detailed algorithm of the proposed recovery-type estimator-enhanced PINN (R-PINN).
 Section 4 presents our results using R-PINN on four representative singular problems, illustrating its performance and providing a comparative analysis.
@@
 The input vector is denoted as $\bm{x}\in\mathbb{R}^{N_{0}}$, and the output vector of the $k$-th layer is expressed as $\mathcal{N}^{k}(\bm{x})$, especially, $\mathcal{N}^{0}(\bm{x})=\bm{x}$.
-The activation function $\Phi$, applied during the feed-forward transition from the $(k-1)$-th layer to the $k$-th layer (for $1\leq k\leq L$), introduces non-linearity into the network, thereby enhancing its learning capacity.
+The activation function $\Phi$, applied during the feed-forward transition from the (k1)(k-1)-th layer to the $k$-th layer (for $1\leq k\leq L$), introduces non-linearity into the network, thereby enhancing its learning capacity.
 Consequently, a feed-forward neural network can be mathematically defined as
-$\mathcal{N}^{k}(\bm{x})=\Phi(\bm{W}^{k}\mathcal{N}^{k-1}(\bm{x})+\bm{b}^{k}),\quad 1\leq k\leq L.$ (2.1)
-Letting $\tilde{\bm{\Theta}}=\{\bm{W}^{k},\bm{b}^{k}\}\in\mathcal{V}$ be all parameters of the neural network, including weights and biases, where $\mathcal{V}$ denotes the parameter space, we can finally express the neural network as
-$\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}}),$
-where $\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ emphasizes the dependence of the output layer $\mathcal{N}^{L}(\bm{x})$ on $\tilde{\bm{\Theta}}$.
+$\mathcal{N}^{k}(\bm{x})=\Phi(\bm{W}^{k}\mathcal{N}^{k-1}(\bm{x})+\bm{b}^{k}),\quad 1\leq k\leq L.$
+Letting $\bm{\Theta}=\{\bm{W}^{k},\bm{b}^{k}\}\in\mathcal{V}$ be all parameters of the neural network, including weights and biases, where $\mathcal{V}$ denotes the parameter space, we can finally express the neural network as
+$\hat{u}_{\bm{\Theta}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\bm{\Theta}),$
+where $\mathcal{N}^{L}(\bm{x};\bm{\Theta})$ emphasizes the dependence of the output layer $\mathcal{N}^{L}(\bm{x})$ on $\bm{\Theta}$.
 Initial parameters are typically generated according to a specified probability distribution.
 In the context of solving PDEs with PINNs, we consider a PDE defined over the domain $\Omega\in\mathbb{R}^{d}$ with parameters $\bm{\lambda}$,
-$f(\bm{x};u(\bm{x}))=f\left(\bm{x};\frac{\partial u}{\partial x_{1}},\cdots,\frac{\partial u}{\partial x_{d}};\frac{\partial^{2}u}{\partial x_{1}^{2}},\cdots,\frac{\partial^{2}u}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)=0,\quad\bm{x}=(x_{1},x_{2},\cdots,x_{d})\in\Omega,$ (2.2)
+$f(\bm{x};u(\bm{x}))=f\left(\bm{x};\frac{\partial u}{\partial x_{1}},\cdots,\frac{\partial u}{\partial x_{d}};\frac{\partial^{2}u}{\partial x_{1}^{2}},\cdots,\frac{\partial^{2}u}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)=0,\quad\bm{x}=(x_{1},x_{2},\cdots,x_{d})\in\Omega,$
 subject to the boundary condition:
@@
 where $u(\bm{x})$ represents the value of the true solution at point $\bm{x}$.
-In solving the forward problem, where the parameter $\bm{\lambda}$ in the equation is known, PINNs approximate the true solution $u$ using a neural network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})$ [1].
-During this process, the neural network is iteratively trained to update its parameters $\tilde{\bm{\Theta}}$ in order to minimize the loss function, defined as,
-$\mathcal{L}(\mathcal{T};\tilde{\bm{\Theta}})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}}),$ (2.3)
+In solving the forward problem, where the parameter $\bm{\lambda}$ in the equation is known, PINNs approximate the true solution $u$ using a neural network $\hat{u}_{\bm{\Theta}}(\bm{x})$ [26].
+During this process, the neural network is iteratively trained to update its parameters $\bm{\Theta}$ in order to minimize the loss function, defined as,
+$\mathcal{L}(\mathcal{T};\bm{\Theta})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\bm{\Theta})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\bm{\Theta}),$ (2.1)
 where
-$\begin{split}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}})=\frac{1}{\lvert\mathcal{T}_{f}\rvert}\sum_{\bm{x}\in\mathcal{T}_{f}}\bigg{|}&f\left(\bm{x};\frac{\partial\hat{u}}{\partial x_{1}},\cdots,\frac{\partial\hat{u}}{\partial x_{d}};\frac{\partial^{2}\hat{u}}{\partial x_{1}\partial x_{1}},\cdots,\frac{\partial^{2}\hat{u}}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)\bigg{|}^{2},\\ &\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}})=\frac{1}{\lvert\mathcal{T}_{b}\rvert}\sum_{\bm{x}\in\mathcal{T}_{b}}\lvert\mathcal{B}(\hat{u},\bm{x})\rvert^{2},\end{split}$ (2.4)
+$\begin{split}\mathcal{L}_{f}(\mathcal{T}_{f};\bm{\Theta})=\frac{1}{\lvert\mathcal{T}_{f}\rvert}\sum_{\bm{x}\in\mathcal{T}_{f}}\bigg|&f\left(\bm{x};\frac{\partial\hat{u}_{\bm{\Theta}}}{\partial x_{1}},\cdots,\frac{\partial\hat{u}_{\bm{\Theta}}}{\partial x_{d}};\frac{\partial^{2}\hat{u}_{\bm{\Theta}}}{\partial x_{1}\partial x_{1}},\cdots,\frac{\partial^{2}\hat{u}_{\bm{\Theta}}}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)\bigg|^{2},\\ &\mathcal{L}_{b}(\mathcal{T}_{b};\bm{\Theta})=\frac{1}{\lvert\mathcal{T}_{b}\rvert}\sum_{\bm{x}\in\mathcal{T}_{b}}\lvert\mathcal{B}(\hat{u}_{\bm{\Theta}},\bm{x})\rvert^{2},\end{split}$
 and $w_{f}$ and $w_{b}$ are weighting factors.
@@
 To solve the inverse problem with PINNs, the loss function not only accounts for the deviations dictated by physical laws but also incorporates an additional term to capture the residual between observed solution $u(\bm{x})$ and predicted solution $\hat{u}(\bm{x})$ over a set of collocation points $\mathcal{T}_{i}$:
-$\mathcal{L}_{i}(\mathcal{T}_{i};\tilde{\bm{\Theta}},\bm{\lambda})=\frac{1}{|\mathcal{T}_{i}|}\sum_{\bm{x}\in\mathcal{T}_{i}}|\hat{u}(\bm{x})-u(\bm{x})|^{2}.$ (2.5)
+$\mathcal{L}_{i}(\mathcal{T}_{i};\bm{\Theta},\bm{\lambda})=\frac{1}{|\mathcal{T}_{i}|}\sum_{\bm{x}\in\mathcal{T}_{i}}|\hat{u}_{\bm{\Theta}}(\bm{x})-u(\bm{x})|^{2}.$
 As a result, the total loss function for the inverse problem can be expressed as:
-$\mathcal{L}(\mathcal{T};\tilde{\bm{\Theta}},\bm{\lambda})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}},\bm{\lambda})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}},\bm{\lambda})+w_{i}\mathcal{L}_{i}(\mathcal{T}_{i};\tilde{\bm{\Theta}},\bm{\lambda}),$ (2.6)
-with the weight of the additional term $w_{i}$. Then the unknown parameters $\bm{\lambda}$ can be trained with the neural network parameters $\tilde{\bm{\Theta}}$ by the optimization process.
+$\mathcal{L}(\mathcal{T};\bm{\Theta},\bm{\lambda})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\bm{\Theta},\bm{\lambda})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\bm{\Theta},\bm{\lambda})+w_{i}\mathcal{L}_{i}(\mathcal{T}_{i};\bm{\Theta},\bm{\lambda}),$ (2.2)
+with the weight of the additional term $w_{i}$. Then the unknown parameters $\bm{\lambda}$ can be trained with the neural network parameters $\bm{\Theta}$ by the optimization process.
 Note that in equations (2.1) and (2.2), it is crucial to position the set of collocation points within the domain of interest.
@@
 2.2 Sampling methods for PINNs
-Typically, the sampling strategy in PINNs involves selecting a set of collocation points by uniform sampling prior to training the neural network, with the set remaining fixed throughout the training process.
-Commonly employed uniform sampling methods include equispaced uniform sampling and uniformly random sampling, while other uniform sampling techniques are also available [1, 4, 30].
-In Wu.
-et al.’s work [19], six uniform sampling methods were systematically reviewed, including equispaced uniform grid (Grid), uniformly random sampling (Random), Latin hypercube sampling (LHS) [31, 32], Halton sequence (Halton) [33], Hammersley sequence (Hammersley) [34], and Sobol sequence (Sobol) [35].
+Typically, the sampling strategy in PINNs involves selecting a set of collocation points by sampling prior to training the neural network, with the set remaining fixed throughout the training process.
+Common strategies for uniformly sampling collocation points include equispaced grid sampling and uniformly random sampling.
+Several other quasi-uniform sampling techniques have also been proposed in the literature [26, 24, 11].
+In Wu et al.’s work [35], six uniform sampling methods were systematically reviewed, including equispaced uniform grid (Grid), uniformly random sampling (Random), Latin hypercube sampling (LHS) [21, 30], Halton sequence (Halton) [13], Hammersley sequence (Hammersley) [14], and Sobol sequence (Sobol) [29].
 Their comparative experiments demonstrated that the Sobol method generally provides superior performance.
-The Sobel sequence uses a base of two to form successively finer uniform partitions of the unit interval and then reorder the coordinates in each dimension.
+The Sobol sequence uses a base of two to form successively finer uniform partitions of the unit interval and then reorders the coordinates in each dimension.
 We show the difference over three different methods in Fig.
 1.
-Fig. 1: $900$ uniformly collocation points generated on $[0,1]^{2}$ with three different strategies.
-In many cases, accuracy of numerical solution can be enough through uniform sampling strategies.
-However, for certain numerical problems characterized by large local gradients and highly distinctive solution features, such methods may fail to adequately capture local variations.
+Fig. 1: $900$ uniform collocation points generated on $0,1]^{2}$ with three different strategies.
+In many cases, accuracy of numerical solution can be good enough through uniform sampling strategies.
+However, for certain numerical problems characterized by large local gradients and highly distinctive solution features, such methods may fail to adequately capture the local variations.
 Several adaptive methods have been proposed to overcome these limitations in recent years, primarily through modifications to the sampling strategies for the collocation points in PINNs.
-These adaptive sampling approaches can be broadly categorized into two types [19]: the Residual-based adaptive refinement (RAR) [18] and the Residual-based adaptive distribution (RAD) [21].
-RAR targets regions with large residuals, where new adaptive collocation points are introduced at regular intervals during the training process to reduce local errors (see Algorithm 1).
-Algorithm 1 Residual-based adaptive refinement (RAR)
-1:  Sample background collocation points $\mathcal{T}$.
-2:  Train the neural network with $\mathcal{T}$ for a certain number of epoches.
-3:  while not reach the end do
-4:     Sample a set of adaptive collocation points $\mathcal{S}_{0}$.
-5:     Compute the PDE residuals of adaptive collocation points $\mathcal{S}_{0}$.
-6:     $S\leftarrow m$ points with largest residuals from $\mathcal{S}_{0}$.
-7:     $\mathcal{T}\leftarrow\mathcal{T}\cup\mathcal{S}$.
-8:     Train the neural network for some epoches.
-9:  end while
-10:  Output the result.
-While RAR only focuses on the area with large residuals and ignore the other place, RAD utilizes the probabilty density function $p(\bm{x})$ that has strong relationship with residual to resample all the collocation points, which can take the area both with large and small residual into care (see Algorithm 2).
-For example, we can take $p(\bm{x})$ proportional to the PDE residual $\epsilon(\bm{x})=|f(\bm{x};\hat{u}(\bm{x})|$,
-[…]
-More forms of $p(x)$ can be used here, such as it in paper[19],
-$p(\bm{x})\propto\frac{\epsilon^{k}(\bm{x})}{\mathbb{E}[\epsilon^{k}(\bm{x})]}+c,$ (2.8)
-where $k\geq 0,c\geq 0$ are two hyperparameters.
-Algorithm 2 Residual-based adaptive distribution (RAD)
-1:  Sample background collocation points $\mathcal{T}$.
-2:  Train the neural network with $\mathcal{T}$ for a certain number of epoches.
-3:  while not reach the end do
-4:     $\mathcal{T}\leftarrow$ resampled collocation points according to the probability density function $p(\bm{x})$.
-5:     Train the neural network for some epoches.
-6:  end while
-7:  Output the result.
+These adaptive sampling approaches can be broadly classified into two categories:
+  • The first category incrementally adds newly selected points based on an adaptive criterion, meaning that previously selected points are retained and the number of sampling points increases over time.
+Representative methods in this category include the residual-based adaptive refinement (RAR) [18], which identifies regions with large residuals and adds new points periodically, and the self-adaptive importance sampling (SAIS) [9], which constructs a probability density function (PDF) based on the residual to sample new points and adds them to the existing dataset.
+  • The second category, in contrast, maintains a fixed total number of sampling points by periodically redistributing them based on an updated sampling criterion.
+A typical example is the residual-based adaptive distribution (RAD) [23], which uses a residual-informed PDF $p(\bm{x})$ to resample all collocation points.
+A recent study [35] shows that methods of the second category often achieve better performance.
+Therefore, in this work, we adopt the strategy of shifting all adaptively distributed points periodically.
+The adaptive sampling methods discussed above are all based on the residual-type error estimator.
+However, another practical alternative, the recovery-type error estimator, has not yet been explored in PINNs-based methods.
+By relying on reconstructed solution fields rather than direct residuals, recovery-based estimators provide a different perspective on error characterization, which can lead to more robust and informative sampling criteria.
+This motivates us to develop an adaptive PINN framework based on recovery-type error estimation, along with a corresponding sampling strategy.
 3 Recovery-type estimator enhanced PINN
-In this section, we present an adaptive PINNs algorithm developed by our team, which utilizes the recovery-type a-posteriori error estimator [27, 28, 29], a widely used technique in adaptive finite element methods.
+In this section, we present an adaptive PINN algorithm, which utilizes the recovery-type a-posteriori error estimator [39, 40, 41], a widely used technique in adaptive finite element methods.
 The concept of adaptive finite element method, first introduced by I.
@@
 in 1978 [3], has since seen significant advancements and remains integral to practical scientific computations.
-This approach extends classical finite element methods by incorporating a-posteriori error estimation and automated mesh refinement to iteratively improve the accuracy of approximate solutions, which can achieve the target accuracy with reduced computational effort, offering distinct advantages over traditional finite element methods, particularly when addressing problems involving singular solutions.
+This approach extends classical finite element methods by incorporating a-posteriori error estimation and automated mesh refinement to iteratively improve the accuracy of approximated solutions, which can achieve the target accuracy with reduced computational effort, offering distinct advantages over traditional finite element methods, particularly when addressing problems involving singular solutions.
 The fundamental concept of the adaptive finite element method lies in the use of a-posteriori error estimator to assess the numerical solution error within each mesh element, thereby guiding the mesh refinement and coarsening process.
 This adaptive mesh strategy shares a notable similarity with adaptive PINNs, where the distribution of collocation points is adjusted based on the error distribution.
-In adaptive finite element methods, commonly employed a-posteriori error estimators can be categorized into two main types: the first is based on residual type [25, 36], where the error is estimated by computing the discrepancy between the solution and the residual of the governing equation; the second type, recovery type, proposed by the Zienkiewicz and Zhu [27, 28, 29], is a gradient reconstruction-based method that provides more accurate error estimates by reconstructing the solution gradients.
-In existing adaptive PINNs algorithms, the equation operator is typically used as the default a-posteriori error estimator, which corresponds to the residual-based error estimator in adaptive finite element methods.
+In adaptive finite element methods, commonly employed a-posteriori error estimators can be categorized into two main types: the first is based on residual type [3, 4], where the error is estimated by computing the discrepancy between the solution and the residual of the governing equation; the second type, recovery type, proposed by Zienkiewicz and Zhu [39, 40, 41], is a gradient reconstruction-based method that provides more accurate error estimates by reconstructing the solution gradients.
+In existing adaptive PINNs algorithms, the PDE residual, computed using automatic differentiation, is typically used as the default a-posteriori error estimator, which corresponds to the residual-based error estimator in adaptive finite element methods.
 Our proposed algorithm seeks to introduce the recovery-type error estimator into the adaptive PINNs framework, with the aim of enriching and extending the theoretical foundation of adaptive PINNs, ultimately providing a more comprehensive approach to error estimation and control.
 3.1 Recovery-type a-posteriori estimator
-In 1987, O.C.
-Zienkiewicz and J.Z.
-Zhu introduced a pioneering a-posteriori error estimator based on gradient recovery for finite element theory, commonly referred to as the Z-Z method [27] or the recovery-type method.
-This approach, notable for its computational simplicity, intuitive understanding, and straightforward implementation within finite element software, quickly gained widespread acceptance within the engineering community.
-Its rapid adoption represented a significant milestone, facilitating the transition from theoretical research to practical application in the domain of finite element a-posteriori error estimation.
-In the context of a specific recovery-type a-posteriori error estimation framework, a typical problem in the field of elasticity is to be considered as an example, which is encapsulated within the following system of equations
-$\bm{L}\bm{u}=\bm{S}^{T}\bm{DS}\bm{u}=\bm{f}\quad\text{in}\ \Omega,$ (3.1)
-which is subject to suitable boundary conditions on the domain’s boundary $\partial\Omega$.
-Here, $\bm{u}$ represents the displacement vector, $\bm{L}$ is the differential operator intrinsic to the theory of elasticity, and the matrix operator $\bm{S}$ defines the strain tensor $\bm{\epsilon}$ by
-$\bm{\epsilon}=\bm{S}\bm{u}.$ (3.2)
-Furthermore, the elasticity matrix $\bm{D}$ is utilized to determine the stress tensor $\bm{\sigma}$, as delineated by the equation:
-$\bm{\sigma}=\bm{D}\bm{\epsilon}.$ (3.3)
-The finite element solution to the aforementioned elasticity problem is expressed in the following form
-$\bm{u}_{h}=\bm{N}\bar{\bm{u}}_{h},$ (3.4)
-where $\bar{\bm{u}}_{h}$ denotes the vector of nodal displacements, and $\bm{N}$ represents the set of approximate polynomial shape functions that interpolate the displacement field within the finite element framework Subsequently, the stress tensor $\bm{\sigma}_{h}$, is computed using the approximate displacement field $\bm{u}_{h}$ with
-$\bm{\sigma}_{h}=\bm{DSu}_{h}.$ (3.5)
-The recovery-type a-posteriori error estimator is characterized by focusing on estimating the numerical error by approximating the gradient recovery of the numerical solution $\bm{\sigma}^{*}$.
-Therefore, the gradient recovery technique plays a key role for the estimator.
-With the development of recovery-type a-posteriori error estimators, there have been many gradient recovery methods, the typical ones are the transformation method and nodal averaging method.
-Fig. 2: The finite element method solution and the smoothened solution of stress for a one-dimensional problem.
-  • 1.
-    The transformation method
-    For this method, it is assumed that the stress
-    $\bm{\sigma}^{*}=\bm{N}\bar{\bm{\sigma}}^{*},$ (3.6)
-    and the method seeks to minimize the discrepancy between the smooth stress $\bm{\sigma}^{*}$ and the stress of the finite element $\bm{\sigma}_{h}$ by enforcing the condition
-[…]
-    By combining equations (3.6) and (3.7), we can derive the following expression for $\bar{\bm{\sigma}}^{*}$
-[…]
-    This method is highly accurate but computationally expensive due to the need to solve the additional system of equations.
-  • 2.
-    The nodal averaging method
-    In this simpler approach, the smooth stress $\bar{\sigma}^{*}$ is obtained by averaging the finite element stresses $\bm{\sigma}_{h,i}$ at the nodes
-    $\bm{\sigma}^{*}=\frac{\left(\sum_{i=1}^{M}\bm{\sigma}_{h,i}\right)}{M}$ (3.9)
-    where $M$ is the number of elements connected with the node $i$.
-    This method is straightforward and computationally efficient, although it may not achieve the same level of accuracy as the transformation method.
-Nevertheless, it performs well in practical applications.
-As for error estimation, two types of errors are usually considered. One is the error in $\bm{\sigma}$ which can be expressed as
-$\lVert\bm{e}_{\sigma}\rVert=\lVert\bm{\sigma}-\bm{\sigma}_{h}\rVert.$ (3.10)
-The other is the recovery-type estimator of the error is similarly defined in the appropriate norm as
-$\lVert\tilde{\bm{e}}_{\sigma}\rVert=\lVert\bm{\sigma}^{*}-\bm{\sigma}_{h}\rVert,$ (3.11)
-and its efficiency is gauged by the efficacy index defined as
-$\theta=\frac{\lVert\tilde{\bm{e}}_{\sigma}\rVert}{\lVert\bm{e}_{\sigma}\rVert}.$ (3.12)
-The confines of this efficacy index are elucidated through the subsequent theorem (as detailed in [29]):
-$1-\frac{\lVert\bm{e}_{\sigma}^{*}\rVert}{\lVert\bm{e}_{\sigma}\rVert}\leq\theta\leq 1+\frac{\lVert\bm{e}_{\sigma}^{*}\rVert}{\lVert\bm{e}_{\sigma}\rVert},$ (3.13)
-with the error in the reconstructed solution characterized as
-$\lVert\bm{e}_{\sigma}^{*}\rVert=\lVert\bm{\sigma}-\bm{\sigma}^{*}\rVert.$ (3.14)
-It is evident that the efficacy index approaches unity when the precision of the reconstructed solution significantly outperforms that of the original finite element solution.
-Should the reconstructed solution exhibit superconvergence, that is,
-$\lVert\bm{e}_{\sigma}^{*}\rVert=C^{*}h^{p+\alpha},$ (3.15)
-where $\alpha$ is a positive constant and $h$ represents the element size, and
-$\lVert\bm{e}_{\sigma}\rVert=Ch^{p},$ (3.16)
-then the bounds for the efficacy index are
-$1-\tilde{C}h^{\alpha}\leq\theta\leq 1+\tilde{C}h^{\alpha},$ (3.17)
-with $\tilde{C}=C^{*}/C$, rendering the error estimator asymptotically precise.
-This estimation underscores the efficiency of the recovery-type a-posteriori estimator.
-Given that the error estimation revolves around stress $\bm{\sigma}$, which is contingent upon the gradients of the solution $\bm{u}$, this class of estimator may also be regarded as a form of gradient recovery.
-3.2 Algorithm of Recovery-type estimator enhanced PINN
+In this section, we provide a brief overview of recovery-type a-posteriori error estimators in the context of finite element methods, along with a discussion of how they work in practice.
+A more detailed theoretical treatment can be found in [36, 31].
+Let $\mathcal{T}_{h}$ be a quasi-uniform triangulation of the domain $\Omega$, and denote by $\mathcal{Z}_{h}$ the set of its vertices.
+For a given node $z\in\mathcal{Z}_{h}$, we define an element patch $\omega$ consisting of all elements that share the vertex $z$.
+We place the origin of a local coordinate system at $z$, and denote by $(x_{j},y_{j})$ the barycenter of each triangle $\tau_{j}\subset\omega$, where $j=1,2,\dots,m$.
+As a model problem, we consider the weak form of a non-self-adjoint problem: Find $u\in H^{1}(\Omega)$ such that
+$B(u,v)=\int_{\Omega}\left[(\mathcal{D}\nabla u+\mathbf{b}u)\cdot\nabla v+cuv\right]\,\mathrm{d}{\bm{x}}=f(v),\quad\forall v\in H^{1}(\Omega),$
+with a Dirichlet boundary condition $u|_{\partial\Omega}=g$, where $\mathcal{D}$ is a symmetric positive-definite $2\times 2$ matrix, $\mathbf{b}\in[L^{\infty}(\Omega)]^{2}$ is a vector-valued function, $f\in(H^{1}(\Omega))^{\prime}$ is a linear functional, and $g\in H^{\frac{1}{2}}(\partial\Omega)$ is a prescribed function defined on the boundary.
+We assume that the coefficients are smooth, and that the bilinear form $B(\cdot,\cdot)$ is continuous and satisfies an inf-sup condition on $H^{1}(\Omega)$, ensuring the well-posedness of the problem.
+The finite element approximation $u_{h}\in\mathcal{V}_{h}\subset H^{1}(\Omega)$ is defined by
+$B(u_{h},v_{h})=f(v_{h}),\quad\forall v_{h}\in\mathcal{V}_{h}.$
+To guarantee a unique discrete solution, we further assume that the inf-sup condition also holds on the discrete space $\mathcal{V}_{h}$.
+To indicate the recovery-type a-posteriori error, we introduce a gradient recovery operator $G$ designed to provide an improved approximation of the true gradient $\nabla u$.
+Ideally, the recovered gradient $Gu_{h}$ satisfies the following inequality:
+$\lVert\nabla u-Gu_{h}\rVert_{0,\Omega}\leq\beta\lVert\nabla u-\nabla u_{h}\rVert_{0,\Omega},\quad\text{with }0\leq\beta<1,$
+which leads to a two-sided bound:
+$(1-\beta)\lVert\nabla u-\nabla u_{h}\rVert_{0,\Omega}\leq\lVert\nabla u_{h}-Gu_{h}\rVert_{0,\Omega}\leq(1+\beta)\lVert\nabla u-\nabla u_{h}\rVert_{0,\Omega}.$
+This justifies the use of $\lVert\nabla u_{h}-Gu_{h}\rVert_{0,\Omega}$ as a recovery-based a-posteriori error estimator—it is fully computable and reliably reflects the true error.
+There are several standard approaches for constructing $Gu_{h}$ at a vertex $z$:
+  1. (a)
+    Weighted averaging: The recovered gradient is constructed as a weighted average of the gradients evaluated at the barycenters of the neighboring triangles:
+    $Gu_{h}(z)=\sum_{j=1}^{m}\frac{|\tau_{j}|}{|\omega|}\nabla u_{h}(x_{j},y_{j}).$
+  2. (b)
+    Local $L^{2}$ projection: For each component $l=1,2$, we seek a linear function $p_{l}\in P_{1}(\omega)$ satisfying
+    $\int_{\omega}\left[p_{l}(x,y)-\partial_{l}u_{h}(x,y)\right]q(x,y)\,\mathrm{d}x\mathrm{d}y=0\quad\forall q\in P_{1}(\omega),$
+    and define $Gu_{h}(z)=(p_{1}(0,0),p_{2}(0,0))$.
+  3. (c)
+    Local least-squares fitting (Zienkiewicz-Zhu method [40]): This method solves a discrete version of the $L^{2}$ projection, by minimizing the difference at the barycenters:
+    $\sum_{j=1}^{m}\left[p_{l}(x_{j},y_{j})-\partial_{l}u_{h}(x_{j},y_{j})\right]q(x_{j},y_{j})=0,\quad\forall q\in P_{1}(\omega),\quad l=1,2.$
+    Again, we define $Gu_{h}(z)=(p_{1}(0,0),p_{2}(0,0))$.
+Subsequently, $Gu_{h}$ can be linearly interpolated from $Gu_{h}(z),\ z\in\mathcal{Z}_{h}$.
+Among these three approaches, the weighted averaging method is particularly popular due to its simplicity and low computational cost.
+The least-squares fitting can be viewed as a discrete analogue of the $L^{2}$ projection.
+Detailed analysis on the convergence and performance of these estimators can be found in [40, 41, 36, 31].
+3.2 Algorithm of the recovery-type estimator enhanced PINN
 In this section, we introduce the algorithm of recovery-type estimator enhanced adaptive PINN (R-PINN).
-The success of the adaptive PINNs algorithm hinges on two crucial components: the posterior error estimation method and the adaptive sampling strategy.
-The posterior error estimation predicts the discrepancy between the numerical and true solutions, while the adaptive sampling strategy adjusts the distribution of collocation points based on this error, allowing for more focused and efficient training in areas that need it most.
-Section 3.1 provides a comprehensive explanation of the posterior error estimation method.
-Therefore, for our R-PINN, the key is how to combine PINNs with this kind of posterior error estimator to realize the adaptive process.
-To solve the equation (2.2), we utilize a shape-regular triangulation $\mathcal{T}_{h}$ formed by a family of partitions of the domain $\Omega$, which allows us to effectively implement the recovery-type a-posteriori estimator.
-We define the operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ as the recovery-type estimator used in our algorithm, employing a nodal averaging method for gradient recovery (other gradient recovery methods can also be utilized).
-The approximated solution $\hat{u}_{\tilde{\Theta}}$ predicts the solution values at all nodes within the mesh $\mathcal{T}_{h}$, which then facilitates gradient recovery and error computation for each element $K\in\mathcal{T}_{h}$ through the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$.
-Fig. 3: The flow chart of R-PINN.
-Another crucial aspect is the selection of collocation points and the adaptive adjustment of their distribution.
-Similar to many adaptive PINNs algorithms, we start with $N_{1}$ background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ for pre-training.
-During the adaptive training phase, we introduce $N_{2}$ adaptive collocation points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ alongside the background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$, updating the adaptive collocation points after each training batch.
-For background collocation points, we utilize a Sobol sequence as described in section 2.2.
-For the adaptive collocation points, we propose a novel approach called the recovery-type estimator based adaptive distribution (RE-AD).
-Unlike common sampling methods, the RE-AD technique is element-based, as the computed a-posteriori errors are distributed across elements.
-This means we only need to specify the number of adaptive collocation points to sample in each element, allowing for random sampling within each element.
-Therefore, it’s vital to allocate adaptive collocation points intelligently to target areas with significant errors without creating redundancy.
-To achieve this, we design an iterative process: we define a tolerance $\epsilon$ such that $\frac{N_{2}}{N_{e}}\leq\epsilon<1$, where $N_{e}$ is the total number of elements.
-Initially, we select $L_{1}=[\epsilon N_{e}]$ elements that exhibit the largest errors in $\mathcal{T}_{h}$, which we refer to as the sequence $\mathcal{S}_{1}$.
-Adaptive collocation points are sampled from $\mathcal{S}_{1}$ in proportion to the errors.
-For instance, for the $i$-th element in $\mathcal{S}_{1}$, the number of adaptive collocation points assigned is $[N_{2}\frac{e_{i}}{\sum_{j}e_{j}}]$.
-Since this process rounds down, some adaptive collocation points may remain unallocated.
-We then repeat this process by selecting $L_{2}=[\epsilon L_{1}]$ of the largest elements and allocating adaptive collocation points according to their errors, continuing until no points are redundant.
-The RE-AD sampling method differs from RAR but is more akin to RAD, as it maintains a fixed total number of adaptive collocation points while only adjusting their distribution during each iteration.
-Algorithm 3 Recovery-type estimator based adaptive distribution (RE-AD)
-1:  Set the number of redundant adaptive collocation points $N_{r}=N_{2}$.
-2:  Denote $L_{0}=N_{e}$.
-3:  $k\leftarrow 1$.
-4:  while $N_{r}$ not 00 do
-5:     Select $L_{k}=[\epsilon L_{k-1}]$ elements with the largest errors as $\mathcal{S}_{k}$.
-6:     Designate adaptive collocation points to $\mathcal{S}_{k}$.
-7:     Update the number of redundant adaptive collocation points $N_{r}$.
-8:     $k\leftarrow k+1$.
-9:  end while
-10:  Output adaptive collocation points $\{\bm{x}_{a}\}_{i=1}^{N_{2}}$.
-To introduce the specific procedures of the recovery-type estimator enhanced adaptive PINN (R-PINN), we utilize the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ along with the recovery-type estimator based adaptive distribution (RE-AD) sampling strategy.
-Initially, we use a set of background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ to train the neural network, aiming to obtain a rough approximation of the true solution, which serves as the pre-training step.
-Following the pre-training phase, an adaptive training iteration process begins.
-In each iteration, the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ is used to compute the errors of the approximated solution $\hat{u}_{\tilde{\Theta}}(\bm{x})$ on the mesh $\mathcal{T}_{h}$.
-Based on these errors, the RE-AD sampling method is then employed to strategically distribute additional adaptive collocation points.
-These adaptive collocation points are combined with the background collocation points to retrain the neural network.
-This iterative process continues until the accuracy improvements become minimal, indicating that the network has converged to a satisfactory approximation of the solution.
-Since the number of collocation points can always stay constant, the algorithm satisfies the characteristics of self-adaptation while ensuring accurate solution.
+The success of the adaptive PINNs algorithm hinges on two crucial components: the a-posteriori error estimation method and the adaptive sampling strategy.
+The a-posteriori error estimation predicts the discrepancy between the numerical and true solutions, while the adaptive sampling strategy adjusts the distribution of collocation points based on this error, allowing for more focused and efficient training on areas that need it the most.
+We begin by discussing a-posteriori error estimation based on the recovery-type estimator.
+The core idea is to utilize the estimator introduced in Section 3.1, particularly the weighted averaging scheme, to assist in evaluating the a-posteriori error.
+The procedure is as follows.
+After training the PINN, we obtain an approximated solution $\hat{u}_{\bm{\Theta}}=\mathcal{N}^{L}(\bm{x};\bm{\Theta})$.
+Given a quasi-uniform triangulation $\mathcal{T}_{h}$ of the domain $\Omega$, we interpolate $\hat{u}_{\bm{\Theta}}$ at mesh vertices to obtain a piecewise linear approximation $\hat{u}_{I}$ such that
+$\hat{u}_{I}(z)=\hat{u}_{\bm{\Theta}}(z),\quad\forall z\in\mathcal{Z}_{h},$
+where $\mathcal{Z}_{h}$ denotes the set of mesh vertices.
+The interpolated solution $\hat{u}_{I}$ serves as an intermediate quantity for both gradient recovery and subsequent a-posteriori error estimation.
+The recovered gradient at a node $z\in\mathcal{Z}_{h}$ is computed using a weighted averaging approach,
+$G\hat{u}_{I}(z)=\sum_{j=1}^{m}\frac{|\tau_{j}|}{|\omega|}\nabla\hat{u}_{I}(x_{j},y_{j}),$
+where $(x_{j},y_{j})$ is the barycenter of each triangle $\tau_{j}\subset\omega$, and $\omega$ is an element patch consisting of all elements that share the vertex $z$.
+The recovered gradient field $G\hat{u}_{I}$ is then obtained by linear interpolation of the nodal values $G\hat{u}_{I}(z)$.
+Using the recovered gradient, the recovery-type a-posteriori error estimator is defined element-wise as
+$\eta_{K}=\lVert\nabla\hat{u}_{I}-G\hat{u}_{I}\rVert_{0,K},\quad K\in\mathcal{T}_{h}.$ (3.1)
+Here, $\eta_{K}$ provides a local indication of the error within element $K$, guiding the identification of regions where additional sampling points are required to enhance the solution accuracy.
+Before presenting the sampling strategy, we outline the overall training strategy. The training process is divided into two phases: a pre-training phase and an adaptive training phase.
+In the pre-training phase, the PINN is trained using only a fixed set of background collocation points, denoted by $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$.
+This phase aims to obtain an initial approximation of the solution.
+The pre-training is followed by an adaptive training phase, during which both background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ and a set of adaptively distributed points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ are employed.
+While the background collocation points remain fixed throughout the entire training process, the adaptively distributed points are dynamically updated in each iteration.
+Specifically, the adaptive phase comprises $M$ iterations.
+At the beginning of each iteration, the recovery-based a-posteriori error estimator $\eta_{K}$ is computed to evaluate the current approximation error over each element $K$.
+Based on these error indicators, new adaptively distributed points are selected in regions with relatively large errors.
+The PINN is then retrained using the union of background and updated adaptively distributed points for a fixed number of epochs before proceeding to the next iteration.
+During the adaptive training process, the sampling strategy plays a key role in determining how the collocation points are selected and updated.
+The background collocation points, denoted by $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$, are generated using a Sobol sequence as described in Section 2.2, and remain fixed throughout the entire training procedure.
+In contrast, the adaptively distributed points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ are updated in each iteration based on the element-wise a-posteriori error indicators $\eta_{K}$}.
+To support the adaptively distributed points selection, we introduce a sampling approach called recovery-type estimator based adaptive distribution (RecAD), which helps allocate adaptively distributed points to regions with larger estimated errors.
+Fig. 2: Flowchart of the R-PINN algorithm.
+Unlike common sampling methods, the RecAD technique is element-based, as the computed a-posteriori errors are distributed across elements.
+This means we only need to specify the number of adaptively distributed points to sample in each element, allowing for random sampling within each element.
+Therefore, a reasonable scheme is required to assign the number of adaptively distributed points across elements based on their estimated a-posteriori errors.
+To this end, we design an iterative procedure.
+We begin by sorting all elements in descending order according to their estimated a-posteriori errors, and denote them by the set $\mathcal{S}_{0}=\{K_{j}\}_{j=1}^{N_{e}}$, where $N_{e}$ is the total number of elements.
+Let $n_{1},n_{2},\dots,n_{N_{e}}$ denote the number of adaptively distributed points assigned to each element, which are initially all set to zero.
+Additionally, we introduce a tolerance parameter $\epsilon\in\left(\frac{1}{N_{e}},1\right)$, representing the proportion of elements to be selected for sampling.
+Algorithm 1 Recovery-type estimator based adaptive distribution (RecAD)
+1: Initialize the number of remaining adaptively distributed points $N_{r}=N_{2}$, the number of remaining elements $m_{0}=N_{e}$, and set the tolerance parameter $\epsilon$.
+2: Sort all elements in descending order according to their estimated a-posteriori errors to obtain the set $\mathcal{S}_{0}=\{K_{j}\}_{j=1}^{N_{e}}$.
+3: Set the initial number of assigned points for each element as $n_{j}=0$, for $j=1,2,\cdots,N_{e}$.
+4: Set iteration counter $k\leftarrow 1$.
+5:while $N_{r}\neq 0$ do
+6:  Select $m_{k}=\left[\epsilon m_{k-1}\right]$ elements with the largest a-posteriori errors from $\mathcal{S}_{k-1}$ as $\mathcal{S}_{k}$.
+7:  Assign the number of adaptively distributed points to elements in $\mathcal{S}_{k}$ proportionally:
+$n_{j}\leftarrow n_{j}+\left[N_{r}\frac{\eta_{K_{j}}}{\sum_{i=1}^{m_{k}}\eta_{K_{i}}}\right],\quad j=1,2,\cdots,m_{k}.$
+8:  Update the number of remaining adaptively distributed points $N_{r}$ by equation (3.2).
+9:  if $m_{k}<\frac{1}{\epsilon}$ then
+10:   Assign the remaining points to the element with the largest error: $n_{1}\leftarrow n_{1}+N_{r}$.
+11:   break
+12:  end if
+13:  $k\leftarrow k+1$.
+14:end while
+15: Randomly sample the adaptively distributed points $\{\bm{x}_{a}\}_{i=1}^{N_{2}}$ with the numbers $n_{1},n_{2},\cdots,n_{N_{e}}$ in each element.
+16:Output: Adaptively distributed points $\{\bm{x}_{a}\}_{i=1}^{N_{2}}$.
+At the first iteration, we select the top $m_{1}=\left[\epsilon N_{e}\right]$ elements with the largest a-posteriori errors from $\mathcal{S}_{0}$, and denote this subset as $\mathcal{S}_{1}=\{K_{j}\}_{j=1}^{m_{1}}$.
+In these selected elements, the number of adaptively distributed points are assigned approximately in proportion to their estimated a-posteriori errors,
+$n_{j}=\left[N_{2}\frac{\eta_{K_{j}}}{\sum_{i=1}^{m_{1}}\eta_{K_{i}}}\right],\quad j=1,2,\cdots,m_{1}.$
+Since rounding is applied to ensure that each $n_{j}$ is an integer, the total number of assigned points is likely to be smaller than the prescribed $N_{2}$.
+We then define the number of remaining points $N_{r}$ as
+$N_{r}=N_{2}-\sum_{j=1}^{N_{e}}n_{j}.$ (3.2)
+In the next iteration, we select a subset of elements from $\mathcal{S}_{1}$, specifically the top $m_{2}=\left[\epsilon m_{1}\right]$ elements with the largest a-posteriori errors, denoted as $\mathcal{S}_{2}=\{K_{j}\}_{j=1}^{m_{2}}$.
+The number of remaining points $N_{r}$ are then redistributed to these elements according to their relative a-posteriori errors
+$n_{j}=n_{j}+\left[N_{r}\frac{\eta_{K_{j}}}{\sum_{i=1}^{m_{2}}\eta_{K_{i}}}\right],\quad j=1,2,\cdots,m_{2}.$
+After each iteration, we update $N_{r}$ using the same equation (3.2).
+The process continues until one of the following stopping criteria is met: (1) the number of adaptively distributed points $N_{2}$ has been fully allocated (namely $N_{r}=0$), or (2) the number of selected elements becomes smaller than $\frac{1}{\epsilon}$, meaning the next subset would be empty.
+In the second case, suppose that after $k$-th iterations we obtain $\mathcal{S}_{k}=\{K_{j}\}_{j=1}^{m_{k}}$ with $N_{r}>0$ and $m_{k}<\frac{1}{\epsilon}$.
+Then, continuing the iteration would result in an empty set $\mathcal{S}_{k+1}$.
+To avoid this, we directly assign the number of remaining points $N_{r}$ to the element with the largest a-posteriori error $K_{1}$, namely $n_{1}=n_{1}+N_{r}$.Now that the number of adaptively distributed points for each element in the mesh $\mathcal{T}_{h}$ has been fully determined, we then randomly sample the corresponding number of points within each element based on the assigned values $n_{1},n_{2},\dots,n_{N_{e}}$, thereby completing the adaptive sampling process.
+The full algorithm for this adaptive distribution scheme RecAD is summarized in Algorithm 1.
+With RecAD in place, all components of the R-PINN framework are complete.
+The entire training procedure of R-PINN is outlined in Algorithm 2.
 Algorithm 2 Recovery-type estimator enhanced PINN (R-PINN)
-0:  The number of adaptive samples and background samples $N_{a}$ and $N_{b}$, the maximum steps number $M$
-1:  Generate $N_{1}$ background samples $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ with Sobol sequence.
-2:  Pre-train the network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ with background samples $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$.
-3:  Set the number of iteration $k\leftarrow 1$ and give the max number of iterations $M$.
+0: Number of adaptively distributed points $N_{a}$, number of background points $N_{b}$, and maximum number of adaptive iterations $M$.
+1: Generate $N_{1}$ background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ with Sobol sequence.
+2: Pre-train the network $\hat{u}_{\bm{\Theta}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\bm{\Theta})$ with background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$.
+3: Set the iteration counter $k\leftarrow 1$.
 4:while $k\leq M$ do
-5:     Generate $N_{2}$ adaptive samples $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ with RE-AD method.
-6:     Retrain the network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ with the dataset $\mathcal{D}=\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}\cup\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$.
-7:  end while
-[…]
+5:  Compute the a-posteriori errors of $\hat{u}_{\bm{\Theta}}(\bm{x})$ with the recovery-type estimator defined in the equation (3.1).
+6:  Generate $N_{2}$ adaptively distributed points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ with RecAD method.
+7:  Retrain the network $\hat{u}_{\bm{\Theta}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\bm{\Theta})$ with the dataset $\mathcal{D}=\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}\cup\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$.
+8:end while
+9:return The approximated solution $\hat{u}_{\bm{\Theta}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\bm{\Theta})$.
 4 Numerical experiments
 In this section, we present numerical experiments demonstrating the effectiveness of our proposed algorithm.
-Our evaluation compares three different kinds of PINNs: the Recovery-type estimator enhanced PINN (R-PINN), Failure-informed PINN (FI-PINN) [20], and standard PINN as a baseline.
+Our evaluation compares three kinds of PINNs: recovery-type estimator enhanced PINN (R-PINN), failure-informed PINN (FI-PINN) [9] and residual-based attention in PINN (RBA-PINN) [1].
+Among these methods, the first two adopt the adaptive sampling strategy, while RBA-PINN employs a strategy of adaptive weighting.
+For the gradient reconstruction in R-PINN, we employ the weighted averaging approach throughout all experiments.
+It is observed that the other two reconstruction methods achieve similarly comparable performance.
 4.1 Experimental setup
-To ensure a fair comparison, we maintain identical parameters and training processes across all methods for the same number of iterations, varying only the number and distribution of adaptive collocation points.
+To ensure a fair comparison, we maintain identical parameters and training processes across all methods for the same number of iterations, varying only the number and distribution of adaptively distributed points.
 The recovery-type a-posteriori estimation is implemented using FEALPy [34].
@@
 The training process comprises two phases: an initial pre-training phase utilizing the L-BFGS optimizer with a learning rate of $0.1$, conducted for a problem-dependent number of epochs, followed by an adaptive training phase consisting of $M$ adaptive iterations.
-Each iteration is constrained to $5000$ epochs while employing the same optimization method.
-For collocation points, we use $N_{1}=2000$ background collocation points and $N_{2}=100$ adaptive collocation points for R-PINN unless specified, while FI-PINN requires $N_{2}=500$ adaptive collocation points to achieve convergence unless specified.
-For the standard PINN, only $N_{1}=2000$ background collocation points are used during training, without incorporating additional adaptive collocation points.
-However, the training process remains the same as that of the two adaptive PINNs, with each adaptive iteration also consisting of 5000 epochs.
+Each iteration is constrained to $5000$ epochs while employing the same optimization method unless otherwise specified.
+Fig. 3: Illustration of a $10\times 10$ regular diagonal triangulation.
+For the choice of parameters, we set $\epsilon=0.02$ for R-PINN.
+The triangulation of the domain adopted in the method is a uniform diagonal mesh composed of right triangles.
+An example with a $10\times 10$ mesh is illustrated in Fig.
+3.
+Unless otherwise specified, we use a $50\times 50$ mesh as the default configuration in the following experiments.
+For collocation points, we use $N_{1}=2000$ background collocation points and $N_{2}=100$ adaptively distributed points for R-PINN unless specified.
+For the RBA-PINN, only $N_{1}=2000$ background collocation points are used during training, without incorporating additional adaptively distributed points unless specified.
+However, the training process remains the same as that of the two adaptive sampling PINNs.
 All methods utilize $200$ boundary collocation points.
-Notably, our approach achieves comparable or superior accuracy despite using fewer total epochs and adaptive collocation points than related studies.
+Notably, our approach achieves comparable or superior accuracy despite using fewer total epochs and adaptively distributed points than related studies.
 The larger $N_{2}$ value for FI-PINN is necessary as it fails to achieve adaptivity with fewer points.
-Furthermore, while our method maintains a fixed number of adaptive collocation points, FI-PINN accumulates points throughout the adaptive iteration process.
+Furthermore, while our method maintains a fixed number of adaptively distributed points, FI-PINN accumulates points throughout the adaptive iteration process.
 To evaluate the accuracy of our method, we employ two error metrics: the relative $L^{2}$ error,
-$\frac{\lVert\hat{u}{\tilde{\Theta}}-u\rVert_{2}}{\lVert u\rVert_{2}},$ (4.1)
-and the $L^{\infty}$ error, defined as $\lVert\hat{u}{\tilde{\Theta}}-u\rVert_{\infty}$.
-Here, $\hat{u}_{\tilde{\Theta}}$ denotes the numerical solution obtained by our method, and $u$ represents the exact solution.
-4.2 Poisson equation
-Consider the following Poisson equation:
-$\begin{split}-\Delta u(x,y)=&f(x,y),\quad(x,y)\ \text{in}\ \Omega,\\ u(x,y)=&g(x,y),\quad(x,y)\ \text{on}\ \partial\Omega,\end{split}$ (4.2)
-where the domain $\Omega$ is $=[0,1]\times[0,1]$, and the true solution is
-$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}.$ (4.3)
-The solution exhibits a sharp peak near the point $(0.5,0.5)$, which is challenging for standard PINN to capture accurately.
-In this experiment, we pre-train the network for $10000$ iterations and set the maximum number of adaptive iterations to $M=5$.
-R-PINN FI-PINN standard PINN
-$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
-$1$ $3.386$ $0.0990$ $100$ $2.717$ $0.6541$ $500$ $0.809$ $0.6270$
-$2$ $0.732$ $0.0282$ $100$ $1.514$ $0.4593$ $1000$ $0.689$ $0.6180$
-$3$ $0.125$ $0.0260$ $100$ $1.157$ $0.3095$ $1500$ $0.650$ $0.5683$
-$4$ $0.285$ $0.0088$ $100$ $0.294$ $0.0686$ $2000$ $0.879$ $0.5689$
-$5$ $0.139$ $0.0086$ $100$ $0.068$ $0.0335$ $2500$ $0.682$ $0.5570$
-Table 1: The relative $L^{2}$ error and $L^{\infty}$ error by two methods solving Poisson equation.
+$\frac{\lVert\hat{u}_{\bm{\Theta}}-u\rVert_{2}}{\lVert u\rVert_{2}},$
+and the $L^{\infty}$ error, defined as $\lVert\hat{u}_{\bm{\Theta}}-u\rVert_{\infty}$.
+Here, $\hat{u}_{\bm{\Theta}}$ denotes the numerical solution obtained by our method, and $u$ represents the exact solution.
+4.2 Poisson’s equation
+Consider the following Poisson’s equation:
+$\begin{split}-\Delta u(x,y)=&f(x,y),\quad(x,y)\ \text{in}\ \Omega,\\ u(x,y)=&g(x,y),\quad(x,y)\ \text{on}\ \partial\Omega,\end{split}$
+where the domain $\Omega=[0,1]\times[0,1]$, and the true solution is
+$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}.$
+The solution exhibits a sharp peak near the point $(0.5,0.5)$, which is challenging for vanilla PINN to capture accurately.
+In this experiment, the network is pre-trained for 15,000 epochs.
+The number of background collocation points is set to $N_{1}=4000$, and the maximum number of adaptive iterations is specified as $M=4$.
+Each adaptive iteration consists of 10,000 epochs instead of 5,000, due to the difficulty of approximating this particular example, which features a solution with a sharp peak.
+R-PINN FI-PINN RBA-PINN
+$N_{\text{iter}}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$
+1 0.1695 0.0259 120 0.4166 0.2862 500 0.3042 0.2040
+2 0.0673 0.0177 120 0.1656 0.0449 1000 0.3360 0.1880
+3 0.0497 0.0103 120 0.1405 0.0160 1500 0.2781 0.1828
+4 0.0277 0.0015 120 0.0530 0.0086 2000 0.2741 0.1791
+Table 1: The relative $L^{2}$ and $L^{\infty}$ error obtained by three different PINNs solving Poisson’s equation.
 R-PINN
@@
 (b) *
-Fig. 4: The distribution of adaptive collocation points in the first $2$ adaptive iterations with two adaptive PINN methods.
+Fig. 4: The distribution of adaptively distributed points in the first 22 adaptive iterations with two adaptive PINN methods.
 Fig.
-4 illustrates the distribution of adaptive collocation points by two adaptive methods in the first two adaptive iterations.
-It is evident that all adaptive collocation points of our model converge to the neighborhood of the point $(0.5,0.5)$, while adaptive collocation points of FI-PINN are more dispersed.
+4 illustrates the distribution of adaptively distributed points by two adaptive sampling methods in the first two adaptive iterations.
+It is evident that all adaptively distributed points of our model converge to the neighborhood of the point $(0.5,0.5)$, while adaptively distributed points of FI-PINN are more dispersed.
 This shows the strong effectiveness of our algorithm in handling problems with significant local variations.
-Table 1 compares the relative $L^{2}$ error and $L^{\infty}$ error among R-PINN, FI-PINN and standard PINN during the adaptive iterations where $N_{\text{iter}}$ represents the number of adaptive iterations.
-The relative $L^{2}$ error for R-PINN reaches $0.139$, and the $L^{\infty}$ error reaches $0.0086$.
-In contrast, the errors obtained by FI-PINN decrease slowly at the beginning, but drop fast later, reaching the relative $L^{2}$ error of $0.068$ and $L^{\infty}$ error of $0.0335$.
+Table 1 compares the relative $L^{2}$ error and $L^{\infty}$ error among R-PINN, FI-PINN and RBA-PINN during the adaptive iterations where $N_{\text{iter}}$ represents the number of adaptive iterations.
+The relative $L^{2}$ error for R-PINN reaches 0.0277, and the $L^{\infty}$ error reaches 0.0015.
+In contrast, the errors obtained by FI-PINN decrease slowly at the beginning, but drop fast later, reaching the relative $L^{2}$ error of 0.0530 and $L^{\infty}$ error of 0.0086.
 It can be observed that R-PINN can converge faster than FI-PINN.
-While, R-PINN can reach a smaller relative $L^{2}$ error in this case with the exponentially accumulated adaptive collocation points, which is one of its advantage.
-As for the normal PINN, the errors obtained show little to no decrease, indicating that the normal PINN fails to converge effectively.
+R-PINN can also reach a smaller relative $L^{2}$ error even in the case where the adaptively distributed points of FI-PINN are continuously accumulated.
+As for RBA-PINN, it appears to have a difficulty in capturing the local features of the solution accurately.
 Similar results are observed in Fig.
-6, where the lowest relative $L^{2}$ error for our algorithm is below $1\times 10^{-1}$.
+6, where the lowest relative $L^{2}$ error for our algorithm is below $5\times 10^{-2}$.
 In the left panel of Fig.
@@
 The middle panel of Fig.
-5 illustrates the cross-section at $y=0.5$.
-The predicted curve using R-PINN fits the true solution very well, while the solution by FI-PINN has a little obvious error at the foot of the peak.
-The final solutions with three methods are plotted in the right of Fig.
+5 illustrates the cross-sections at $y=0.5$.
+Both R-PINN and FI-PINN are able to accurately approximate the peak-shaped solution.
+In contrast, although RBA-PINN identifies the local feature, its accuracy is limited by the insufficient number of collocation points in the corresponding region.
+The final solutions by three different PINNs are plotted in the right of Fig.
 5.
@@
 (b) *
-standard PINN
+RBA-PINN
 (c) *
-Fig. 5: The prediction err (left), the cross section at $y=0.5$ (middle) and the predicted solution of three methods solving Poisson equation.
-Fig. 6: The relative $L^{2}$ errors of three methods in the optimization process of solving Poison equation.
+Fig. 5: The absolute errors across the domain (left), the cross-sections at $y=0.5$ (middle) and the predicted solutions of three adaptive methods solving Poisson’s equation.
+Fig. 6: The relative $L^{2}$ error by three different PINNs in the optimization process of solving Poisson’s equation.
 4.3 Burgers’ equation
 We consider the following Burgers’ equation:
-$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\frac{0.01}{\pi}\frac{\partial^{2}u}{\partial x^{2}},\\ &u(x,0)=-\sin(\pi x),\\ &u(t,-1)=u(t,1)=0\end{split}$ (4.4)
+$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\frac{0.001}{\pi}\frac{\partial^{2}u}{\partial x^{2}},\\ &u(x,0)=-\sin(\pi x),\\ &u(t,-1)=u(t,1)=0\end{split}$
 where the domain is $(x,t)\in[-1,1]\times[0,1]$.
-The solution to this Burgers’ equation exhibits a sharp mutation along $x=0$, making it prone to errors and offsets.
-For this test, we pre-train the network for $5000$ iterations and use $M=4$ steps of adaptive processes.
-R-PINN FI-PINN standard PINN
-$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
-$1$ $0.0086$ $0.0735$ $100$ $0.2278$ $1.8142$ $200$ $0.0136$ $0.1583$
-$2$ $0.0107$ $0.0964$ $100$ $0.0131$ $0.1263$ $400$ $0.0144$ $0.1685$
-$3$ $0.0071$ $0.0673$ $100$ $0.0064$ $0.0576$ $600$ $0.0174$ $0.2099$
-$4$ $0.0059$ $0.0439$ $100$ $0.0089$ $0.0846$ $800$ $0.0197$ $0.2418$
-Table 2: The relative $L^{2}$ error and $L^{\infty}$ error by three methods solving Burgers’ equation.
+The solution to this Burgers’ equation exhibits an extremely sharp mutation along $x=0$, which makes the problem highly susceptible to numerical errors and local deviations.
+In this test, we pre-train the network with $N_{1}=50,000$ background collocation points for $15,000$ iterations and use $M=4$ steps of adaptive processes.
+R-PINN FI-PINN RBA-PINN
+$N_{\text{iter}}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$
+1 0.0166 0.3260 500 0.2106 1.8845 600 0.1692 1.9880
+2 0.0169 0.4087 500 0.1997 1.9430 1200 0.1692 1.9881
+3 0.0122 0.2078 500 0.1926 1.9593 1800 0.1693 1.9881
+4 0.0048 0.0920 500 0.1928 1.9471 2400 0.1693 1.9881
+Table 2: The relative $L^{2}$ and $L^{\infty}$ error obtained by three different PINNs solving Burgers’ equation.
 R-PINN
@@
 (b) *
-Fig. 7: The distribution of adaptive collocation points in the first $2$ adaptive iterations with two adaptive PINN methods solving Burgers’ equation.
-The adaptive collocation points generated during the first two adaptive training processes by two adaptive PINN method are illustrated in Fig.
+Fig. 7: The distribution of adaptively distributed points in the first 22 adaptive iterations with two adaptive PINN methods solving Burgers’ equation.
+The adaptively distributed points generated during the first two adaptive training processes by two adaptive sampling PINN methods are illustrated in Fig.
 7.
-We observe that the adaptive collocation points successfully cluster around $x=0$ with both methods, where the error is largest (as seen in Fig.
+We observe that the adaptively distributed points successfully cluster around $x=0$ with both methods, where the error is largest (as seen in Fig.
 8).
@@
 The exact relative $L^{2}$ error and $L^{\infty}$ error are listed in Table 2.
-As shown, the final relative $L^{2}$ error and $L^{\infty}$ error for R-PINN are $0.0059$ and $0.0439$, respectively, while those for FI-PINN are $0.0089$ and $0.0846$.
-The prediction err of standard PINN keeps at the save level during the training, where relative $L^{2}$ error and $L^{\infty}$ reach $0.0197$ and $0.2418$ in the final.
-Similar to the experiment of Poisson equation, it takes a long iterations for FI-PINN before the error jumps quickly, while R-PINN can reduce the errors at once.
-The errors obtained by R-PINN and FI-PINN are one order of magnitude lower than those obtained by the standard PINN.
-Such behaviour can also be observed in the Fig.
-10, where the relative $L^{2}$ errors are illustrated specifically, showing that the lowest error can reach below $1\times 10^{-2}$.
-We also show the absolute errors across the entire domain and the cross-section at $t=1$ in Fig.
+As shown, the final relative $L^{2}$ error and $L^{\infty}$ error for R-PINN are 0.0048 and 0.0920, respectively, while those for FI-PINN are 0.1928 and 1.9471.
+In contrast, the final relative $L^{2}$ and $L^{\infty}$ errors of RBA-PINN are 0.1693 and 1.9881, respectively.
+Both FI-PINN and RBA-PINN fail to accurately capture the shock, whereas R-PINN exhibits significantly lower errors than the other two methods.
+Such behaviour can also be observed in Fig.
+10, where the relative $L^{2}$ error is illustrated specifically, showing that the lowest error can reach below $1\times 10^{-2}$.
+We also show the absolute errors across the entire domain and the cross-sections at $t=1$ in Fig.
 8.
-The predicted solutions of two adaptive methods are illustrated in Fig.
+The predicted solutions of two adaptive sampling methods are illustrated in Fig.
 9.
@@
 (b) *
-standard PINN
+RBA-PINN
 (c) *
-Fig. 8: The prediction err (left) and the cross section at $t=1$ (right) of three methods solving Burgers’ equation.
+Fig. 8: The absolute errors across the domain (left) and the cross-sections at $t=1$ (right) of three different PINNs solving Burgers’ equation.
 Fig. 9: The predicted solution (R-PINN at the left and FI-PINN at the right).
-Fig. 10: The relative $L^{2}$ errors of three methods in the optimization process of solving Burgers’ equation.
+Fig. 10: The relative $L^{2}$ error by three different PINNs in the optimization process of solving Burgers’ equation.
 4.4 Two-dimensional problem with two peaks
 Considering the following two-dimensional problem
-$\begin{split}-\nabla\cdot[u(x,y)\nabla(x^{2}+y^{2})]+\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in\Omega,\\ u(x,y)=b(x,y),\quad(x,y)\in\partial\Omega,\end{split}$ (4.5)
+$\begin{split}-\nabla\cdot[u(x,y)\nabla(x^{2}+y^{2})]+\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in\Omega,\\ u(x,y)=b(x,y),\quad(x,y)\in\partial\Omega,\end{split}$
 where $\Omega=[-1,1]\times[-1,1]$. The true solution is
-$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}+e^{-1000[(x+0.5)^{2}+(y+0.5)^{2}]},$ (4.6)
-which has two peaks at $(-0.5,-0.5)$ and $(0.5,0.5)$.
+$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}+e^{-1000[(x+0.5)^{2}+(y+0.5)^{2}]},$
+which has two peaks at (0.5,0.5)(-0.5,-0.5) and $(0.5,0.5)$.
 This test mainly can show the ability of R-PINN to deal with multiple peaks.
@@
 Since two peaks problem is challenging for neural network to approximate, we use $5$ hidden layers with $64$ neurons in each layer to enhance its capability.
-Moreover, we use $N_{2}=150$ adaptive collocation points for R-PINN, and $N_{2}=750$ added adaptive collocation points for FI-PINN.
-R-PINN FI-PINN standard PINN
-$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
-$1$ $0.6778$ $0.1228$ $150$ $0.4372$ $0.4845$ $750$ $0.7912$ $0.6651$
-$2$ $0.1202$ $0.0198$ $150$ $0.5220$ $0.3429$ $1500$ $0.8020$ $0.6475$
-$3$ $0.1080$ $0.0259$ $150$ $0.4773$ $0.2233$ $2250$ $0.8012$ $0.6483$
-$4$ $0.0840$ $0.0127$ $150$ $0.3969$ $0.2263$ $3000$ $0.8342$ $0.6625$
-Table 3: The relative $L^{2}$ error and $L^{\infty}$ error by two methods solving two-dimensional problem with two peaks.
+Moreover, we use $N_{2}=150$ adaptively distributed points for R-PINN, and $N_{2}=750$ added adaptively distributed points for FI-PINN.
+R-PINN FI-PINN RBA-PINN
+$N_{\text{iter}}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$
+1 0.6778 0.1228 150 0.4372 0.4845 750 0.4045 0.5299
+2 0.1202 0.0198 150 0.5220 0.3429 1500 0.3764 0.4885
+3 0.1080 0.0259 150 0.4773 0.2233 2250 0.3578 0.4646
+4 0.0840 0.0127 150 0.3969 0.2263 3000 0.3474 0.4507
+Table 3: The relative $L^{2}$ and $L^{\infty}$ error obtained by three different PINNs solving the two-dimensional problem with two peaks.
 R-PINN
@@
 (b) *
-Fig. 11: The distribution of adaptive collocation points in the first two adaptive iterations of two adaptive methods solving the two-dimensional problem with two peaks.
+Fig. 11: The distribution of adaptively distributed points in the first two adaptive iterations of two adaptive sampling methods solving the two-dimensional problem with two peaks.
 R-PINN
@@
 (b) *
-standard PINN
+RBA-PINN
 (c) *
-Fig. 12: The predicted err (left), the cross section at $y=-0.5$ (middle) and the cross section at $y=0.5$ (right) of three methods solving the two-dimensional problem with two peaks.
-Fig. 13: The predicted solution (R-PINN at the left, FI-PINN at the middle and standard PINN at the right) of the two-dimensional problem with two peaks.
-Fig. 14: The relative $L^{2}$ errors of three methods in the optimization process of solving the two-dimensional problem with two peaks.
-We present the distribution of adaptive collocation points of two adaptive methods across the first two adaptive iterations in Fig.
+Fig.
+12: The absolute errors across the domain (left), the cross-sections at y=0.5y=-0.5 (middle) and the cross-sections at $y=0.5$ (right) of three different PINNs solving the two-dimensional problem with two peaks.
+Fig. 13: The predicted solution (R-PINN at the left, FI-PINN at the middle and RBA-PINN at the right) of the two-dimensional problem with two peaks.
+Fig. 14: The relative $L^{2}$ error by three different PINNs in the optimization process of solving the two-dimensional problem with two peaks.
+We present the distribution of adaptively distributed points of two adaptive sampling methods across the first two adaptive iterations in Fig.
 11.
-It is evident that these points converge around the vicinities of $(0.5,0.5)$ and $(-0.5,-0.5)$ with R-PINN, demonstrating the capability of our algorithm to effectively handle problems characterized by multiple regions of abrupt changes.
-However, the distribution of adaptive collocation points with FI-PINN seems not so effective to deal with the two peaks.
+It is evident that these points converge around the vicinities of $(0.5,0.5)$ and (0.5,0.5)(-0.5,-0.5) with R-PINN, demonstrating the capability of our algorithm to effectively handle problems characterized by multiple regions of abrupt changes.
+However, the distribution of adaptively distributed points with FI-PINN seems not so effective to deal with the two peaks.
 The reason of it might be a truncated Gaussian distribution is used, so that problems of more than one regions with large error are difficult to handle.
 This problem is also reflected in the training process.
-Table 3 compares the relative $L^{2}$ and $L^{\infty}$ errors among R-PINN, FI-PINN and standard PINN during the adaptive iterations.
+Table 3 compares the relative $L^{2}$ and $L^{\infty}$ errors among R-PINN, FI-PINN and RBA-PINN during the adaptive iterations.
 For R-PINN, the relative $L^{2}$ error reduces to 0.0840, and the $L^{\infty}$ error decreases to 0.0127.
-In contrast, the relative $L^{2}$ error associated with FI-PINN can only reach $0.3969$ and the $L^{\infty}$ can get to $0.2263$, which is just a little bit lower than those with standard PINN.
+In contrast, the relative $L^{2}$ error associated with FI-PINN can only reach 0.3969 and the $L^{\infty}$ error can get to 0.2263, while RBA-PINN achieves a lower relative $L^{2}$ error 0.3474 but a larger $L^{\infty}$ error 0.4507.
 Similar trends are observed in Fig.
@@
 The left panel illustrates the absolute errors across the entire domain, revealing significantly smaller errors for R-PINN.
-The middle and right panels display cross-sectional views at $y=-0.5$ and $y=0.5$, respectively, showing that the predictions from R-PINN align closely with the exact solution, whereas there are large errors near two peaks with FI-PINN and the standard PINN can only approximate the features.
+The middle and right panels display cross-sectional views at y=0.5y=-0.5 and $y=0.5$, respectively, showing that the predictions from R-PINN align closely with the exact solution, whereas there are large errors near two peaks for FI-PINN and RBA-PINN.
 Finally, the numerical solutions obtained by the three methods are visualized in Fig.
 13.
+R-PINN FI-PINN RBA-PINN
+$N_{\text{iter}}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$
+1 0.0572 0.0761 100 0.1110 0.1590 500 0.1153 0.0458
+2 0.0076 0.0098 100 0.0317 0.0491 1000 0.1140 0.0453
+3 0.0076 0.0099 100 0.0317 0.0490 1500 0.1137 0.0452
+4 0.0076 0.0099 100 0.0270 0.0418 2000 0.1135 0.0452
+Table 4: The relative $L^{2}$ and $L^{\infty}$ error obtained by three different PINNs solving the wave equation.
 4.5 Wave equation
-For this test, we consider a wave equation as follow,
-$\begin{split}&\frac{\partial^{2}u}{\partial t^{2}}-3\frac{\partial^{2}u}{\partial x^{2}}=0,\quad(t,x)\in[0,6]\times[-5,5],\\ &u(0,x)=\frac{1}{\cosh(2x)}-\frac{0.5}{\cosh(2(x-10))}-\frac{0.5}{\cosh(2(x+10))},\\ &\frac{\partial u}{\partial t}(0,x)=0,\\ &u(t,-5)=u(t,5)=0,\end{split}$ (4.7)
+For this test, we consider a wave equation as follows,
+$\begin{split}&\frac{\partial^{2}u}{\partial t^{2}}-3\frac{\partial^{2}u}{\partial x^{2}}=0,\quad(t,x)\in[0,6]\times[-5,5],\\ &u(0,x)=\frac{1}{\cosh(2x)}-\frac{0.5}{\cosh(2(x-10))}-\frac{0.5}{\cosh(2(x+10))},\\ &\frac{\partial u}{\partial t}(0,x)=0,\\ &u(t,-5)=u(t,5)=0,\end{split}$
 where the true solution is
-$u(t,x)=\frac{0.5}{\cosh(2(x-\sqrt{3}t))}-\frac{0.5}{\cosh(2(x-10+\sqrt{3}t))}+\frac{0.5}{\cosh(2(x+\sqrt{3}t))}-\frac{0.5}{\cosh(2(x+10-\sqrt{3}t))}.$ (4.8)
+$u(t,x)=\frac{0.5}{\cosh(2(x-\sqrt{3}t))}-\frac{0.5}{\cosh(2(x-10+\sqrt{3}t))}+\frac{0.5}{\cosh(2(x+\sqrt{3}t))}-\frac{0.5}{\cosh(2(x+10-\sqrt{3}t))}.$
 Fig. 15: The plot of the exact solution.
@@
 (b) *
-standard PINN
-(c) *
-Fig. 16: The predicted solution (left), predicted error (middle) and the distribution of adaptive collocation points (right) of three methods solving the wave equation.
-Fig. 17: The relative $L^{2}$ errors of three methods in the optimization process of solving the Poison equation the wave equation.
-R-PINN FI-PINN standard PINN
-$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $L^{2}$ Relative $L^{2}$ $L^{\infty}$
-$1$ $0.0572$ $0.0761$ $100$ $0.1110$ $0.1590$ $500$ $0.3448$ $0.4554$
-$2$ $0.0076$ $0.0098$ $100$ $0.0317$ $0.0491$ $1000$ $0.3158$ $0.4214$
-Table 4: The relative $L^{2}$ error and $L^{\infty}$ error by three methods solving the wave equation.
-In this test, the R-PINN method quickly reaches its performance limit.
-Therefore, we adopt a single adaptive iteration process ($M=2$) and pre-train the network for 5000 iterations.
-We show the relative $L^{2}$ errors and $L^{\infty}$ errors in the Table 4.
+Fig. 16: The predicted solutions (left) and the distribution of adaptively distributed points (right) of two adaptive sampling methods solving the wave equation.
+Fig. 17: Comparison of absolute errors distributions across the domain for three different PINNs: R-PINN (left), FI-PINN (middle), and RBA-PINN (right).
+Fig. 18: The relative $L^{2}$ error by three different PINNs in the optimization process of solving the wave equation.
+In this test, we adopt the adaptive iteration process with the number of $M=4$ and pre-train the network for 5000 iterations.
+We show the relative $L^{2}$ error and $L^{\infty}$ error in the Table 4.
 The relative $L^{2}$ error and $L^{\infty}$ error can reach 0.0076 and 0.0098 with R-PINN, which are obviously lower than the other two methods.
-The detailed decreasing process is convenient to be seen in Fig.
-17 The exact solution is depicted in Fig.
+The decreasing trend of the relative $L^{2}$ error during training can be clearly observed in Fig.
+18.
+The exact solution is depicted in Fig.
 15.
 The left panel of Fig.
-16 compares the predicted solutions of the three methods.
-The middle panel of Fig.
-16 illustrates the absolute errors across the domain, showing that the errors associated with R-PINN are smaller than those with FI-PINN and standard PINN.
+16 compares the predicted solutions of the two adaptive sampling methods.
+Fig.
+17 illustrates the absolute errors across the domain, showing that the errors associated with R-PINN are smaller than those with FI-PINN and RBA-PINN.
 Finally, the right panel of Fig.
-16 displays the adaptive collocation points for two adaptive methods and random points for the standard PINN.
-From the right panel, we can see that the same problem occurs for FI-PINN when dealing with multiple large errors region, so it can’t handle well.
+16 displays the adaptively distributed points for two adaptive sampling methods.
+From the right panel, it can be observed that FI-PINN encounters similar difficulties when resolving multiple regions with large errors, indicating its limited effectiveness in such scenarios, where R-PINN demonstrates better performance.
+Fig. 19: Comparison of the relative $L^{2}$ error (left) and $L^{\infty}$ error (right) under varying mesh sizes for different values of $N_{2}$.
+4.6 Tests on the effects of sampling and mesh
+For R-PINN, two critical factors influencing accuracy are the number of collocation points and the mesh size.
+To evaluate their effects, we conduct experiments on the Poisson’s equation presented in Section 4.2 and examine how these two aspects affect the performance of R-PINN.
+The collocation points consist of two components: the background collocation points and the adaptively distributed points.
+We begin by analyzing the combined influence of $N_{2}$ and the mesh size.
+Specifically, we fix $N_{1}=8000$ and vary the mesh size among $10\times 10$, $30\times 30$, $50\times 50$, and $70\times 70$, while also testing $N_{2}=25,\ 50,\ 100,\ 200$.
+The finest mesh used is $70\times 70$, which already generates 9800 elements.
+A finer mesh would result in only a small number of elements containing background collocation points, thus undermining the effectiveness of the method.
+We first examine the effect of mesh size, illustrated in Fig.
+3.
+A decreasing trend in the $L^{\infty}$ error is clearly observed with finer meshes, whereas the relative $L^{2}$ error does not follow a consistent trend.
+This is reasonable because the mesh is primarily used to guide a-posteriori error estimation and sampling, rather than to enrich the representational capacity of the solution space.
+Next, we analyze the impact of $N_{2}$, which can also be seen in Fig.
+19 by comparing four different curves in each subfigure.
+Both the relative $L^{2}$ error and the $L^{\infty}$ error generally decrease as $N_{2}$ increases, with the exception of a significant increase in relative $L^{2}$ error when $N_{2}$ increases from 100 to 200.
+This anomaly results from an imbalance between the background collocation points and adaptively distributed points, where excessive adaptively distributed points destroy the intended sampling strategy.
+Fig. 20: Comparison of the relative $L^{2}$ error (left) and $L^{\infty}$ error (right) under varying values of $N_{1}$.
+Finally, we examine the effect of the number of background collocation points $N_{1}$.
+With the mesh size fixed at $50\times 50$ and $N_{2}=100$, we evaluate the error under different values of $N_{1}$, as shown in Fig.
+20.
+It is clearly observed that both the relative $L^{2}$ and $L^{\infty}$ errors decrease with increasing $N_{1}$, indicating that more background points contribute to a more accurate solution.
 5 Conclusion
 In this paper, we present recovery-type a-posteriori estimator enhanced adaptive PINN that combines the PINN with a recovery type estimator which is popularly used in the adaptive finite element method.
-Our innovation lies in the use of the recovery-type a-posteriori error estimator, which is different from the residual of PINNs, and the design of a tailored sampling method R-AD.
-We test the performance of R-PINN in several PDE problems with sharp or singular solutions, which normal PINN struggles to handle effectively.
-The results demonstrate that R-PINN achieves significant improvements over normal PINN.
+Our innovation lies in the use of the recovery-type a-posteriori error estimator, which is different from the residual of PINNs, and the design of a tailored sampling method RecAD.
+We compare the performance of the proposed R-PINN with FI-PINN and RBA-PINN on several problems featuring sharp gradients or singularities, which vanilla PINNs struggle to resolve effectively.
+The results demonstrate that R-PINN achieves significant improvements.
 More studies can be done to analyze the performance of R-PINN for more complex problems and we shall focus on and analyze these issues in our future work.
+6 Acknowledgment
+JJ was partially supported by China Natural National Science Foundation (No.
+22341302), the Fundamental Research Funds for the Central Universities, JLU (93Z172023Z05), and the Key Laboratory of Symbolic Computation and Knowledge Engineering of Ministry of Education of China housed at Jilin University.
+YL acknowledged the support of NSF-DMS (No.
+2208499).
+CZ was partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (No.
+XDB0640000) and the National Natural Science Foundation of China (No.
+12571445).
+The authors wish to thank the anonymous reviewers for their constructive comments and valuable suggestions, which significantly improved the quality of this manuscript.